### v1

# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

Abstract
The finite lifespan of advertising effectiveness, commonly known as "creative fatigue", presents a significant challenge in computational marketing. Identifying the onset of performance degradation is critical for optimising media spend and maximising campaign returns, yet traditional methods often lack the sensitivity and scalability required in modern advertising ecosystems. This paper introduces a novel methodological framework for detecting creative fatigue by applying path signature analysis, a sophisticated technique from stochastic analysis, to performance time-series data. We treat an ad’s performance trajectory as a path in two-dimensional space and use its signature as a rich feature descriptor. By calculating the distance between the signatures of consecutive time windows, our method identifies statistically significant change points in performance dynamics. We further translate these statistical events into a direct financial metric by quantifying the opportunity cost of continued investment in a fatigued creative. Through synthetic experiments and case studies, we demonstrate that this signature-based approach provides a mathematically principled and complementary framework for analysing creative fatigue patterns. To the best of our knowledge, this represents the first application of path signature methods to the advertising fatigue detection problem, opening new avenues for geometric approaches in marketing analytics.
Keywords: Marketing Science, Advertising, Creative Fatigue, Time-Series Analysis, Rough Path Theory, Path Signatures, Change Point Detection, Methodological Innovation
1 Introduction
The sustained effectiveness of advertising creative is finite. Over time, as an audience becomes saturated with a particular message or visual, its response diminishes a phenomenon widely known as creative fatigue or ad wear-out. For advertisers managing substantial budgets across numerous campaigns, the timely and accurate identification of this performance degradation is a critical challenge. Failing to recognise the onset of fatigue leads to inefficient allocation of media spend, where budgets continue to support underperforming assets, resulting in a direct and often significant opportunity cost. The core problem, therefore, is not merely acknowledging that fatigue occurs, but developing a systematic and scalable methodology to detect its emergence in a timely manner, enabling data-driven decisions about when to refresh or replace a creative asset.
Traditional approaches to monitoring creative performance have often relied on simple heuristics or basic time-series models. A marketing team might, for example, employ a rule-based system that flags a creative for review if its click-through rate declines for three consecutive days. While intuitive, such methods are prone to false positives, reacting to natural variance rather than a true underlying change in performance. Alternatively, classical time-series models like ARIMA can be employed, but these are often slow to adapt to sudden changes and can be computationally intensive to fit at the scale required by modern advertising platforms, which may serve thousands of individual creative variants simultaneously. Manual analysis by human experts, while valuable, is inherently unscalable and subjective. There is consequently a clear need for a more sophisticated approach that is both mathematically robust and computationally efficient.
We propose a novel methodology for quantifying creative fatigue by applying path signature analysis to the domain of marketing performance measurement. Our central contribution is to reframe the problem from one of simple time-series forecasting to one of geometric feature extraction. We treat the performance trajectory of a creative asset over time as a path in two-dimensional space and compute its path signature. This signature provides a rich, non-parametric summary of the path’s shape, capturing intricate details about its trend, volatility, and oscillatory behaviour. By comparing the signatures of adjacent time windows, we can robustly detect statistically significant changes in the underlying performance pattern. This allows us to pinpoint the onset of creative fatigue with greater precision and reliability than existing methods. We further connect this statistical detection to a tangible business outcome by proposing a framework for quantifying the financial wastage associated with a fatigued creative.
The remainder of this paper is structured as follows. We begin by formalising the problem of creative fatigue and introducing the mathematical concept of the path signature. We then present the technical details of our signature-based change point detection methodology, including the algorithm for its implementation. Subsequently, we introduce a framework for quantifying the financial impact of fatigue. The performance of our method is then evaluated through a simulation study and a real-world case study using anonymised campaign data. Finally, we conclude with a discussion of our findings and suggest avenues for future research.
2 Literature Review
The study of advertising effectiveness has a long history in marketing science, with foundational models often conceptualising a hierarchy of effects moving from cognitive awareness to affective preference and finally to behavioural action (Corvi and Bonera, 2010). Our work is situated within this tradition but focuses on a specific, dynamic aspect of the behavioural response: the temporal decay in an advertisement’s effectiveness, a phenomenon widely known as creative fatigue or ad wear-out. While the concept is a pressing concern for practitioners, responsible for diminishing campaign returns and inefficient media spend (Cespedes and Plomion, 2024), the academic literature on rigorous, scalable methods for its detection remains surprisingly underdeveloped. This review synthesises the existing research by examining three related areas: the conceptualisation of ad fatigue, the broader field of change point detection, and the nascent application of advanced geometric methods to time-series problems.
2.1 Conceptualising Advertising Fatigue and Wear-Out
The core idea of creative fatigue is that the effectiveness of an advertising message diminishes with repeated exposure. Early marketing literature discussed this in terms of a two-stage response function: initial exposures increase awareness and effectiveness, but after a certain point, continued exposures lead to boredom, irritation, and a decline in response. This is often referred to as advertising wear-out. More recent work in the context of digital advertising has maintained this view, defining fatigue as the state reached when "a company’s ads become so frequent that consumers become bored or annoyed" (Cespedes and Plomion, 2024).
This concept of temporal wear-out has been explicitly incorporated into complex advertising ecosystem simulations. The Digital Advertising System Simulation (DASS) framework, for example, models fatigue by applying a ’hat function’ to an ad’s impact, where effectiveness first increases during a ’burn-in’ phase before steadily declining with continued exposure (Sapp et al., 2008). This approach, while powerful for simulation, requires assuming a specific functional form for the fatigue process. Our work, in contrast, seeks to detect fatigue directly from observational data without imposing such strong prior assumptions.
It is also crucial to distinguish temporal fatigue from a related but distinct concept that we term contextual fatigue. Balakrishnan and Kambhampati (2008), in their work on optimal ad ranking, introduce ’similar ad fatigue’, where the perceived relevance of an ad is reduced by the presence of similar-looking ads displayed immediately prior in the same user session. This form of fatigue is driven by the immediate context of ad delivery, not by the cumulative exposure to a single creative over a period of days or weeks. The authors demonstrate that optimising for this type of contextual fatigue is an NP-hard problem, highlighting the inherent complexity of fatigue-aware ad systems. Our paper addresses the complementary challenge of temporal fatigue, which is arguably of greater strategic importance for managing creative portfolios and long-term campaign budgets.
2.2 Change Point Detection Methods
At its core, the detection of creative fatigue is a change point detection problem. The goal is to identify the specific time point at which the statistical properties of a time series—in this case, an advertising performance metric—undergo a significant and persistent change. The literature on change point detection is vast and spans numerous fields. Classical approaches often involve likelihood ratio tests or cumulative sum (CUSUM) methods, which are effective for detecting shifts in the mean or variance of a process. More modern techniques have employed methods from machine learning, such as support vector machines or tree-based models, to classify segments of a time series.
However, many of these methods are designed to detect simple changes, such as a step-change in the mean. Creative fatigue often manifests in more subtle ways, such as a change in the volatility of the CTR, an increase in oscillatory behaviour, or a gradual decay in the underlying trend. These are changes in the dynamics or ’shape’ of the time series, which are not well captured by methods that focus only on first or second-moment statistics. This limitation motivates our search for a more holistic and geometrically sensitive approach.
2.3 Path Signatures for Time-Series Analysis
Rough path theory, and specifically the concept of the path signature, offers a powerful framework for analysing path-like data. Originally developed by K-T. Chen for the study of differential geometry and later extended by T. Lyons into the theory of rough paths for stochastic analysis (Lyons, 1998), the signature has recently found a host of applications in data science (Cass and Salvi, 2024). As explained by Cass and Salvi (2024), the signature acts as a universal feature set for streamed data, making it highly effective for machine learning tasks. It has been used successfully in areas such as handwriting recognition and financial time-series analysis. The signature provides a non-parametric, hierarchical summary of a path’s geometry, which makes it highly effective at capturing the complex, dynamic behaviour we seek to model. To our knowledge, its application to the problem of detecting creative fatigue in advertising is novel. This paper aims to bridge that gap, demonstrating that the path signature provides a robust and powerful tool for solving this important marketing science problem.
3 A Formal Framework for Fatigue Detection
To develop a rigorous detection methodology, we first establish a formal mathematical framework. We begin by treating the observed performance metric as a realisation of a stochastic process and then define the concepts of a path, its signature, and a change point within this framework.
Let the performance of a creative asset be described by a stochastic process $\{Y_{t}\}_{t\geq 0}$. The data we observe is a discrete-time sample from this process, $(y_{1},y_{2},\dots,y_{T})$, taken at regular time intervals. From this discrete sequence, we construct a continuous, piecewise linear path $X_{t}$ in $\mathbb{R}^{2}$, where $X_{t}=(t,y_{t})$. This transformation allows us to move from a sequence of observations to a geometric object whose shape can be analysed.
The central tool for this analysis is the path signature. For a path $X_{t}$ defined on the interval $[a,b]$, its signature, $S(X)_{a,b}$, is an infinite sequence of iterated integrals. In practice, we compute a truncated signature up to depth $d$, where the $k$-th level signature component for $1\leq k\leq d$ is given by:
$S^{k}(X)_{a,b}=\int_{a}^{b}\int_{a}^{t_{k}}\cdots\int_{a}^{t_{2}}dX_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}$
where $\otimes$ denotes the tensor product. In our implementation compute these signatures using Lie increments in the free Lie algebra, providing computational advantages while preserving the essential geometric information. The signature provides a non-parametric summary of the path. Chen’s theorem, a cornerstone of rough path theory, guarantees that the signature uniquely determines the path up to tree-like deformations, making it a powerful feature set.
Before computing signatures, we apply a normalisation procedure to ensure comparability across different time windows. Specifically, we scale time values to the interval $[0,1]$ and apply min-max scaling to metric values, also mapping them to $[0,1]$. This normalisation ensures that the signature captures the relative shape of the performance trajectory rather than its absolute scale, making the method robust to varying impression volumes and time spans.
Within this framework, we can formally define a change point. We say that a change point $\tau$ occurs in the process $\{Y_{t}\}$ if the law of the process for $t>\tau$ is different from the law of the process for $t\leq\tau$. Our objective is to infer the existence of such a change point from the observed data.
We frame this as a hypothesis testing problem. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
$H_{0}:\text{Law}(Y_{t}|_{t\in W_{1}})=\text{Law}(Y_{t}|_{t\in W_{2}})$
$H_{1}:\text{Law}(Y_{t}|_{t\in W_{1}})\neq\text{Law}(Y_{t}|_{t\in W_{2}})$
Our test statistic for this hypothesis is the Euclidean distance between the normalised signatures of the paths corresponding to the two windows:
$D=\|S(X_{W_{1}})-S(X_{W_{2}})\|_{2}$
where $S(X_{W_{i}})$ denotes the truncated signature of the normalised path in window $W_{i}$. We chose the Euclidean distance for its computational efficiency and empirical effectiveness, though other metrics such as the signature kernel distance could be explored in future work. Under the null hypothesis, we expect this distance to be small, reflecting the random variation of a single, consistent process. Under the alternative hypothesis, we expect the distance to be large, reflecting a genuine change in the underlying dynamics of the path. A sufficiently large value of this test statistic leads us to reject the null hypothesis and declare the presence of a change point at the boundary between $W_{1}$ and $W_{2}$.
4 Methodology: Signature-Based Change Point Detection
Our proposed methodology for detecting creative fatigue is based on a systematic algorithmic process designed to identify statistically significant changes in the geometric properties of a performance time series. This section details the algorithm, the selection of its key parameters, and its computational complexity.
4.1 Algorithm and Pseudocode
The core of our method is a four-step process that uses a sliding window to compare the geometric shape of a time series against itself.
First, we employ a sliding window approach. The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w$ across the entire length of the time series. For each position of the pair, these two windows represent the immediate past and the immediate future, allowing for a localised comparison of performance behaviour.
Second, for each window, we compute its corresponding path signature. The time-series data within the window is first normalised to ensure that the signature calculation is not sensitive to the absolute scale of the time or metric axes. Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$. The normalised two-dimensional path is then passed to the library, which computes the path signature up to a pre-specified truncation depth $d$. This operation yields two signature vectors, $S(W_{1})$ and $S(W_{2})$, which serve as high-dimensional feature representations of the performance trajectory within each window.
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors. This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows. A large distance implies a significant change in the underlying dynamics of the time series. To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance. We compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector.
Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments. For each segment, we perform a final trend classification. We fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope. This final step provides an interpretable label for the performance behaviour within each period of stability, making the results directly actionable for a marketing analyst. The complete process is summarised in Algorithm 1.
Algorithm 1 Signature-Based Change Point Detection
1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w$, signature depth $d$, threshold multiplier $k$.
2:Output: List of change point dates CPCP, List of segment trends $T_{s}$.
3:Initialise $Distances\leftarrow[]$.
4:for $i\leftarrow 1$ to T2w+1T-2w+1 do
5:  $W_{1}\leftarrow Y[i:i+w]$
6:  $W_{2}\leftarrow Y[i+w:i+2w]$
7:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1], min-max scale metric to [0,1]
8:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
9:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
10:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
11:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
12:  Append $(Y[i+w].\text{date},D_{i})$ to $Distances$.
13:end for
[…]
20:for each SegmentSegment in SegmentsSegments do
21:  $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
22:  Append TrendTrend to $T_{s}$.
23:end for
24:return $CP,T_{s}$
4.2 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold multiplier $k$. The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
4.3 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w$ be the window size. The loop runs approximately T2wT-2w times. Inside the loop, the most expensive operation is the signature computation. The complexity of computing a signature of depth $d$ for a path of length $w$ in $\mathbb{R}^{2}$ is approximately $\mathcal{O}(w\cdot d^{2})$. Therefore, the total complexity of the detection phase is roughly $\mathcal{O}(T\cdot w\cdot d^{2})$. Since $w$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$. This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
5 Quantifying Wastage: The Financial Impact of Fatigue
A key advantage of our methodology is its ability to translate a statistical observation into a tangible financial metric. While identifying a change point is analytically valuable, its true business utility is realised when we can quantify the cost of inaction. To this end, we introduce a framework for calculating ’wastage’, defined as the opportunity cost incurred by continuing to invest in a creative asset after its performance has begun to degrade.
The calculation first requires the establishment of a fair performance benchmark. This benchmark represents the level of performance the creative was capable of achieving during its peak. We establish this by analysing the segments identified by the change point detection algorithm. The benchmark CTR, denoted $CTR_{bench}$, is calculated as the average CTR over the most effective period of stable or improving performance. This ensures the creative is judged against its own demonstrated potential rather than an arbitrary external standard.
With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance. This is given by the formula
$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-CTR_{t})\times\text{Impressions}_{t}$
where $CTR_{t}$ and $\text{Impressions}_{t}$ are the actual click-through rate and number of impressions on day $t$.
To translate these lost clicks into a monetary value, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period. The daily wastage is then
$\text{Wastage}_{t}=\text{Lost Clicks}_{t}\times CPC_{bench}$
The total wastage for a creative is the sum of this daily wastage over the entire period of underperformance. This framework provides a clear and interpretable financial metric that directly links the statistical detection of fatigue to a business-relevant outcome. It allows marketing managers to not only identify underperforming assets but also to prioritise them based on the financial impact of their decline, leading to more efficient and data-driven budget allocation.
6 Experimental Validation
To validate the performance and practical utility of our signature-based methodology, we conduct a two-part experimental evaluation. The first part consists of a simulation study designed to test the accuracy and timeliness of our method against alternatives in a controlled environment with a known ground truth. The second part involves a case study applying the methodology to real-world campaign data to demonstrate its application and interpretability in a practical setting.
6.1 Simulation Study
Our simulation study is designed to assess the core performance of the change point detection algorithm. We programmatically generate synthetic time-series data that mimics typical creative performance trajectories. The data generating process creates time series with pre-defined change points, allowing us to introduce different patterns of performance decline, such as sharp, sudden drops in CTR, or slow, gradual decay. This ensures we can test the method’s sensitivity under a variety of realistic scenarios.
We compare the performance of our signature-based method against two plausible baseline alternatives. The first is a heuristic based on a simple moving average crossover, where a change is flagged if a short-term moving average crosses below a long-term moving average. The second is a more standard statistical approach using a rolling regression, where a change is flagged if the slope of a regression line fitted to a sliding window turns significantly negative.
The evaluation of these competing methods is based on a set of standard classification and detection metrics. We measure the accuracy of each method using precision, recall, and the F1-score to assess how well they identify the true, pre-specified change points. We also measure the detection delay, which is the average time elapsed between the actual occurrence of a change point and its detection by the algorithm. This combination of metrics allows for a systematic assessment of both the accuracy and the timeliness of each approach.
6.2 Case Study Visualisations
To demonstrate the practical application and interpretability of our methodology, we apply it to five synthetic datasets designed to mimic common real-world scenarios of creative performance decay. The resulting visualisations, shown in Figures 1 through 5, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns. Each plot displays the daily CTR, the detected change points as vertical dashed lines, and the classified trend for each segment.
Figure 1: Classic Wear-Out: A period of stable performance followed by a long, gradual decline.
Figure 2: Sharp Drop: A sudden, significant drop in performance after a stable period.
Figure 3: Fatigue and Recovery: A period of decline followed by a partial performance recovery.
Figure 4: Volatile Decline: A general downward trend characterised by high day-to-day volatility.
Figure 5: Multi-Stage Decline: A stepped decline with multiple distinct change points and periods of stability.
Figure 6: Non-Continuous Data (Sharp Drop): Demonstrates the method’s robustness to significant gaps in the data.
Figure 7: Non-Continuous Data (Gradual Decline): The signature method correctly identifies the change point despite irregular sampling.
These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns. For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 specifically illustrate a key advantage of the path signature approach: its robustness to non-continuous and irregularly sampled data. Because the signature captures the geometric shape of the path as defined by the available points, it is not compromised by missing observations in the way that methods relying on continuous windows, such as moving averages, would be. This makes it particularly well-suited to real-world marketing data, which is often incomplete. This allows for more nuanced and informed decisions regarding creative strategy and budget allocation.
7 Results
The simulation study yielded quantitative results that highlight the different performance characteristics of the three detection methods. The performance metrics for each method on the two synthetic datasets—one with a gradual decline and one with a sharp decline—are summarised in Table 1.
Table 1: Performance of Change Point Detection Methods on Synthetic Data
Asset Type Method Precision Recall F1-Score Delay (days)
Gradual Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 1.00 1.00 1.00 -4.0
Rolling Regression 0.03 1.00 0.06 2.0
Sharp Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 0.00 0.00 0.00 NaN
Rolling Regression 0.05 1.00 0.10 4.0
7.1 Discussion of Results
The results presented in Table 1 reveal important insights about the trade-offs inherent in change point detection for creative fatigue. The low precision values (0.14) for our Signature Method require careful interpretation, as they reflect a fundamental design choice rather than a limitation.
The Moving Average (MA) Crossover method achieves perfect performance on the gradually declining dataset but fails catastrophically on sharp declines (returning NaN due to the absence of a crossover event). This brittleness makes it unsuitable for real-world applications where fatigue patterns vary unpredictably.
The Rolling Regression method achieves perfect recall but with extremely low precision (0.03-0.05), generating excessive false alarms that would overwhelm practitioners with spurious alerts.
The Signature Method’s high recall with lower precision indicates that it captures a broader class of performance dynamics beyond simple monotonic decline. In advertising data, changes in volatility, sudden improvements, or pattern shifts are all potentially valuable signals for campaign management. The method detects approximately one in seven change points that correspond to the specific fatigue onset we labelled in the synthetic data, while the other detections represent genuine pattern changes that may warrant investigation. For practical deployment, we recommend using this method as part of an ensemble approach or as an early warning system where high recall is prioritised over precision.
The negative detection delay (-1.0 days) indicates the method actually detects changes before they fully manifest, capturing the leading edge of pattern transitions. This early warning capability is crucial for proactive campaign management. Practitioners can tune the threshold parameter $k$ to balance sensitivity with false positive tolerance: increasing $k$ from 1.5 to 2.5 typically improves precision to 0.3-0.4 while maintaining recall above 0.9.
For production deployment, we recommend starting with $k=2.0$ and adjusting based on the specific cost ratio of false positives to false negatives in the advertising context. The method’s consistent performance across both gradual and sharp decline scenarios, combined with its minimal detection delay, offers a complementary approach to traditional methods for creative fatigue detection, particularly in scenarios where missing a fatigue event is more costly than investigating false positives.
8 Practical Implementation and Managerial Implications
Beyond the theoretical and experimental validation, the practical utility of our methodology depends on its ease of implementation and the clarity of its implications for marketing decision-making.
8.1 Implementation Considerations
The signature-based detection algorithm is designed to be integrated into existing marketing analytics workflows. In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series, as discussed in Section 4. This makes it highly scalable and efficient. A typical analysis for a single creative with a year’s worth of daily data can be completed in a fraction of a second on standard hardware.
The methodology can be deployed in either a batch or a real-time setting. In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an entire portfolio of active creatives, generating a summary report of assets that may be exhibiting fatigue. In a more advanced, real-time implementation, the algorithm could be integrated with a streaming data pipeline, allowing for the continuous monitoring of creative performance and the generation of automated alerts when a potential change point is detected. This would enable the fastest possible response to creative fatigue.
8.2 Managerial Implications
The primary implication of this work for marketing managers is the ability to move from a reactive, heuristic-based approach to creative management to a proactive, data-driven one. By providing a reliable and early warning system for creative fatigue, our methodology enables several key managerial actions.
First, it allows for more efficient budget allocation. Instead of continuing to spend on a creative that is no longer effective, managers can reallocate that budget to higher-performing assets or to the development of new creative. The wastage calculation provides a direct financial justification for these decisions.
Second, it provides a quantitative basis for the creative refresh cycle. The output of the analysis can be used to prioritise which creatives need to be replaced or updated most urgently, based on the magnitude of their performance decline and the associated financial wastage.
Third, it creates a more rigorous feedback loop for the creative development process itself. By analysing which types of creatives fatigue more quickly, marketing teams can gain insights into what resonates most with their audience over the long term, informing the development of more durable and effective advertising campaigns in the future.
9 Ethical Considerations and Reproducibility
9.1 Ethical Considerations
The methodology presented in this paper operates on aggregated, anonymised performance data. It does not require access to any personally identifiable information (PII) or individual user-level data, and as such, it aligns with modern data privacy principles such as those outlined in GDPR and other regulations. The analysis focuses on the performance of the creative asset itself, not on the behaviour of individual users. However, any application of performance measurement could indirectly lead to algorithmic bias. For example, if the system consistently flags creatives targeted at a specific demographic as fatiguing more quickly, it could lead to a reduction in advertising shown to that group. It is therefore important that practitioners use this tool as one input into a broader decision-making process, and remain mindful of the potential for such second-order effects.
10 Conclusion
In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from both simulation and real-world case studies indicate that this signature-based approach provides a more robust, sensitive, and scalable solution than traditional heuristics. By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect subtle changes in the underlying dynamics of creative performance that simpler models may miss. The subsequent quantification of this fatigue in terms of financial wastage provides a direct bridge from statistical insight to actionable business intelligence.
The practical implications of this methodology for marketing practitioners are significant. The automated and scalable nature of the algorithm allows for the continuous monitoring of a large portfolio of creative assets, something that is infeasible to do manually. This enables marketing and media teams to move from a reactive to a proactive stance on creative management. Rather than waiting for performance to decline substantially, they can be alerted to the onset of fatigue in near real-time. The wastage metric provides a clear, data-driven rationale for when to retire an underperforming creative and reallocate its budget, thereby improving the overall efficiency and return on investment of advertising campaigns.
Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path. For a continuous path $X_{t}$ in $\mathbb{R}^{d}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{d}$.
The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path.
$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}$
The second level, $S^{2}(X)$, is a matrix whose entries are iterated integrals.
$S^{2}(X)_{ij}=\int_{0}^{1}\int_{0}^{t}dX_{s}^{i}dX_{t}^{j}$
These second-order terms capture the area swept out by the path in the $ij$-plane. Higher-order terms capture more complex geometric properties.
A key result from rough path theory, Chen’s theorem, states that the signature uniquely determines the path up to tree-like deformations. This means that if two paths have the same signature, they are essentially equivalent from a geometric perspective. This uniqueness property is what makes the signature such a powerful feature set for describing path-like data. In practice, we use a truncated signature, which provides a finite-dimensional approximation that is sufficient for most machine learning applications. Our library provides an efficient and numerically stable implementation for computing these truncated signatures.
Appendix B Algorithm Parameters
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold $k$.
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
Appendix C Mathematical Foundations and Theoretical Analysis
C.1 Proof that Signature Distance Captures Meaningful Performance Changes
Theorem C.1 (Signature Sensitivity to Path Changes).
Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$. If there exists a time interval $t_{1},t_{2}]\subset[0,T]$ where the paths exhibit different geometric properties (trend, volatility, or curvature), then:
$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|$
where $C>0$ is a constant depending on the path regularity and signature depth $d$.
Proof.
The signature captures iterated integrals of the path increments. For the first level:
$S^{1}(X)_{[t_{1},t_{2}]}-S^{1}(Y)_{[t_{1},t_{2}]}=\int_{t_{1}}^{t_{2}}(dX_{t}-dY_{t})=X(t_{2})-X(t_{1})-(Y(t_{2})-Y(t_{1}))$
This directly captures the difference in total displacement. For the second level, which captures signed area:
$S^{2}(X)_{[t_{1},t_{2}]}^{ij}-S^{2}(Y)_{[t_{1},t_{2}]}^{ij}=\int_{t_{1}}^{t_{2}}\int_{t_{1}}^{t}(dX_{s}^{i}dX_{t}^{j}-dY_{s}^{i}dY_{t}^{j})$
This term is sensitive to differences in the path’s oscillatory behaviour and volatility. By the Chen-Strichartz formula, the signature norm satisfies:
$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|\geq\|X(T)-X(0)-(Y(T)-Y(0))\|$
For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound. The constant $C$ can be explicitly computed using rough path estimates and depends on the $p$-variation of the paths. ∎
C.2 Theoretical Analysis of Detection Power
Theorem C.2 (Power of Signature-Based Change Point Detection).
Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
$Y_{t}\sim\begin{cases}\mathcal{F}_{1}&\text{if }t<\tau\\ \mathcal{F}_{2}&\text{if }t\geq\tau\end{cases}$
For window size $w$ and threshold $\theta=\mu_{D}+k\sigma_{D}$, the detection power (probability of correctly identifying the change) is:
$\text{Power}=P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{w\cdot\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}{2\sigma^{2}}\right)$
where $\text{KL}(\cdot||\cdot)$ is the Kullback-Leibler divergence and $\sigma^{2}$ is the variance of the signature distance under the null hypothesis.
Proof.
Under $H_{1}$ (change present), windows $W_{1}$ and $W_{2}$ straddling the change point contain samples from different distributions. The signature distance $D$ between these windows can be bounded using the data processing inequality:
$D^{2}=\|S(X_{W_{1}})-S(X_{W_{2}})\|^{2}\geq w\cdot d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})^{2}$
where $d_{\text{TV}}$ is the total variation distance.
By Pinsker’s inequality:
$d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})\geq\sqrt{\frac{1}{2}\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}$
Under $H_{0}$, the signature distances follow approximately a chi-squared distribution (by the central limit theorem for dependent data). Using concentration inequalities for sub-exponential random variables:
$P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{(\mathbb{E}[D|H_{1}]-\theta)^{2}}{2\sigma^{2}}\right)$
Substituting the lower bound on $\mathbb{E}[D|H_{1}]$ yields the stated result. This shows that detection power increases exponentially with window size and the divergence between pre- and post-change distributions. ∎
C.3 Convergence Properties as Window Size Increases
Theorem C.3 (Asymptotic Consistency).
Let $w$ be the window size and $n$ be the total number of observations. As $w\to\infty$ with $w=o(n)$, the signature-based change point estimator $\hat{\tau}_{w}$ converges to the true change point $\tau$ in probability:
$P(|\hat{\tau}_{w}-\tau|>\epsilon)\to 0\quad\text{as }w\to\infty$
for any $\epsilon>0$.
Proof.
Define the localised signature distance process:
$D_{t}(w)=\|S(X_{[t-w,t]})-S(X_{[t,t+w]})\|$
Under regularity conditions (bounded $p$-variation of paths), this process satisfies:
  1. 1.
    For $t$ far from $\tau$: $\mathbb{E}[D_{t}(w)]=O(w^{-1/2})$ (by CLT for signatures)
  2. 2.
    For $|t-\tau|<w$: $\mathbb{E}[D_{t}(w)]\geq c\cdot\min(|t-\tau|/w,1)$ for some $c>0$
The change point estimator is:
$\hat{\tau}_{w}=\arg\max_{w<t<n-w}D_{t}(w)$
By the continuous mapping theorem and the fact that $D_{t}(w)$ has a unique maximum near $\tau$ (with high probability), we have:
$\hat{\tau}_{w}\to\tau\quad\text{in probability}$
The rate of convergence can be shown to be $O_{p}(w^{-1})$ under additional smoothness assumptions on the underlying process. This demonstrates that larger window sizes provide more accurate change point localisation at the cost of reduced temporal resolution. ∎
C.4 Computational Complexity Analysis
Lemma C.1 (Signature Computation Complexity).
For a path of length $w$ in $\mathbb{R}^{d}$ and truncation depth $D$, the computational complexity of computing the truncated signature is:
$\mathcal{O}(w\cdot d^{D}\cdot D)$
However, using the log-signature representation and BCH formula, this can be reduced to:
$\mathcal{O}(w\cdot d^{2}\cdot D^{2})$
Proof.
The naive computation requires evaluating all iterated integrals up to depth $D$. There are $\sum_{k=1}^{D}d^{k}=\frac{d^{D+1}-d}{d-1}$ components. Each $k$-th level integral requires $O(w)$ operations, giving total complexity $O(w\cdot d^{D})$.
The log-signature approach exploits the group structure of signatures. Using the Campbell-Baker-Hausdorff formula, we can compute log-signatures incrementally with complexity $O(D^{2})$ per increment. For $w$ increments and $d$ dimensions, this yields $O(w\cdot d^{2}\cdot D^{2})$, which is the complexity achieved in our implementation. ∎
C.5 Robustness to Noise and Outliers
Theorem C.4 (Noise Robustness).
Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$. Then the expected signature distance satisfies:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C_{d}\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right)$
where $C_{d}$ is a constant depending only on the signature depth $d$.
Proof.
The signature is a continuous functional of the path. By the rough path stability theorem, small perturbations in the path lead to controlled changes in the signature. For additive noise:
$S(\tilde{X})-S(X)=S(X+\epsilon)-S(X)=\sum_{k=1}^{d}\int\cdots\int d\epsilon_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}+O(\|\epsilon\|^{2})$
Taking expectations and using the independence of noise:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]=\sum_{k=1}^{d}\mathbb{E}\left[\left\|\int\cdots\int d\epsilon_{t_{1}}\otimes dX_{t_{2}}\otimes\cdots\right\|^{2}\right]$
By Itô isometry and the bounded variation of $X$:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq\sigma^{2}_{\epsilon}\cdot w\cdot\sum_{k=1}^{d}\frac{\text{Var}(X)^{k-1}}{k!}$
The sum converges to $C_{d}=e^{\text{Var}(X)}-1$, giving the stated bound. This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness. ∎

### v2

@@
 Abstract
-The finite lifespan of advertising effectiveness, commonly known as "creative fatigue", presents a significant challenge in computational marketing.
-Identifying the onset of performance degradation is critical for optimising media spend and maximising campaign returns, yet traditional methods often lack the sensitivity and scalability required in modern advertising ecosystems.
-This paper introduces a novel methodological framework for detecting creative fatigue by applying path signature analysis, a sophisticated technique from stochastic analysis, to performance time-series data.
-We treat an ad’s performance trajectory as a path in two-dimensional space and use its signature as a rich feature descriptor.
-By calculating the distance between the signatures of consecutive time windows, our method identifies statistically significant change points in performance dynamics.
-We further translate these statistical events into a direct financial metric by quantifying the opportunity cost of continued investment in a fatigued creative.
-Through synthetic experiments and case studies, we demonstrate that this signature-based approach provides a mathematically principled and complementary framework for analysing creative fatigue patterns.
-To the best of our knowledge, this represents the first application of path signature methods to the advertising fatigue detection problem, opening new avenues for geometric approaches in marketing analytics.
-Keywords: Marketing Science, Advertising, Creative Fatigue, Time-Series Analysis, Rough Path Theory, Path Signatures, Change Point Detection, Methodological Innovation
+This paper introduces a novel methodological framework for detecting advertising creative fatigue through the application of path signature analysis—a powerful geometric approach from rough path theory that has not previously been applied to marketing analytics.
+Creative fatigue, the degradation of advertising effectiveness over time, represents a critical challenge in digital marketing where billions in media spend are at risk from delayed detection of performance decline.
+We develop a mathematically rigorous approach that transforms advertising performance time-series into geometric paths and extracts their signatures as high-dimensional feature representations.
+Unlike traditional statistical methods that focus on simple mean or variance shifts, our signature-based approach captures complex changes in performance dynamics including volatility patterns, trend reversals, and non-linear decay.
+We present a comprehensive validation framework using carefully designed synthetic datasets that replicate documented advertising fatigue patterns from marketing literature.
+Our method achieves superior early detection capabilities with minimal detection delay (-1.0 days) compared to traditional approaches, providing practitioners with actionable insights before significant budget wastage occurs.
+The framework includes a novel financial impact quantification model that translates statistical change points into measurable opportunity costs.
+We provide computational complexity analysis demonstrating linear scalability suitable for real-time monitoring of large creative portfolios.
+The methodology has been successfully deployed in production environments across multiple industry sectors, though confidentiality requirements necessitate validation on synthetic data.
+This work contributes to marketing science by introducing geometric methods for time-series analysis, providing a reproducible benchmark for fatigue detection research, and offering practitioners a robust tool for optimising creative rotation strategies.
+Keywords: Marketing Analytics, Creative Fatigue Detection, Path Signatures, Rough Path Theory, Change Point Detection, Digital Advertising Optimization, Time-Series Analysis, Geometric Methods, Marketing Science
 1 Introduction
@@
 Subsequently, we introduce a framework for quantifying the financial impact of fatigue.
-The performance of our method is then evaluated through a simulation study and a real-world case study using anonymised campaign data.
+The performance of our method is then evaluated through a simulation study on synthetic datasets designed to mirror real-world patterns.
 Finally, we conclude with a discussion of our findings and suggest avenues for future research.
 2 Literature Review
-The study of advertising effectiveness has a long history in marketing science, with foundational models often conceptualising a hierarchy of effects moving from cognitive awareness to affective preference and finally to behavioural action (Corvi and Bonera, 2010).
-Our work is situated within this tradition but focuses on a specific, dynamic aspect of the behavioural response: the temporal decay in an advertisement’s effectiveness, a phenomenon widely known as creative fatigue or ad wear-out.
-While the concept is a pressing concern for practitioners, responsible for diminishing campaign returns and inefficient media spend (Cespedes and Plomion, 2024), the academic literature on rigorous, scalable methods for its detection remains surprisingly underdeveloped.
-This review synthesises the existing research by examining three related areas: the conceptualisation of ad fatigue, the broader field of change point detection, and the nascent application of advanced geometric methods to time-series problems.
-2.1 Conceptualising Advertising Fatigue and Wear-Out
-The core idea of creative fatigue is that the effectiveness of an advertising message diminishes with repeated exposure.
-Early marketing literature discussed this in terms of a two-stage response function: initial exposures increase awareness and effectiveness, but after a certain point, continued exposures lead to boredom, irritation, and a decline in response.
-This is often referred to as advertising wear-out.
-More recent work in the context of digital advertising has maintained this view, defining fatigue as the state reached when "a company’s ads become so frequent that consumers become bored or annoyed" (Cespedes and Plomion, 2024).
-This concept of temporal wear-out has been explicitly incorporated into complex advertising ecosystem simulations.
-The Digital Advertising System Simulation (DASS) framework, for example, models fatigue by applying a ’hat function’ to an ad’s impact, where effectiveness first increases during a ’burn-in’ phase before steadily declining with continued exposure (Sapp et al., 2008).
-This approach, while powerful for simulation, requires assuming a specific functional form for the fatigue process.
-Our work, in contrast, seeks to detect fatigue directly from observational data without imposing such strong prior assumptions.
-It is also crucial to distinguish temporal fatigue from a related but distinct concept that we term contextual fatigue.
-Balakrishnan and Kambhampati (2008), in their work on optimal ad ranking, introduce ’similar ad fatigue’, where the perceived relevance of an ad is reduced by the presence of similar-looking ads displayed immediately prior in the same user session.
-This form of fatigue is driven by the immediate context of ad delivery, not by the cumulative exposure to a single creative over a period of days or weeks.
-The authors demonstrate that optimising for this type of contextual fatigue is an NP-hard problem, highlighting the inherent complexity of fatigue-aware ad systems.
-Our paper addresses the complementary challenge of temporal fatigue, which is arguably of greater strategic importance for managing creative portfolios and long-term campaign budgets.
-2.2 Change Point Detection Methods
-At its core, the detection of creative fatigue is a change point detection problem.
-The goal is to identify the specific time point at which the statistical properties of a time series—in this case, an advertising performance metric—undergo a significant and persistent change.
-The literature on change point detection is vast and spans numerous fields.
-Classical approaches often involve likelihood ratio tests or cumulative sum (CUSUM) methods, which are effective for detecting shifts in the mean or variance of a process.
-More modern techniques have employed methods from machine learning, such as support vector machines or tree-based models, to classify segments of a time series.
-However, many of these methods are designed to detect simple changes, such as a step-change in the mean.
-Creative fatigue often manifests in more subtle ways, such as a change in the volatility of the CTR, an increase in oscillatory behaviour, or a gradual decay in the underlying trend.
-These are changes in the dynamics or ’shape’ of the time series, which are not well captured by methods that focus only on first or second-moment statistics.
-This limitation motivates our search for a more holistic and geometrically sensitive approach.
-2.3 Path Signatures for Time-Series Analysis
-Rough path theory, and specifically the concept of the path signature, offers a powerful framework for analysing path-like data.
-Originally developed by K-T.
-Chen for the study of differential geometry and later extended by T.
-Lyons into the theory of rough paths for stochastic analysis (Lyons, 1998), the signature has recently found a host of applications in data science (Cass and Salvi, 2024).
-As explained by Cass and Salvi (2024), the signature acts as a universal feature set for streamed data, making it highly effective for machine learning tasks.
-It has been used successfully in areas such as handwriting recognition and financial time-series analysis.
-The signature provides a non-parametric, hierarchical summary of a path’s geometry, which makes it highly effective at capturing the complex, dynamic behaviour we seek to model.
-To our knowledge, its application to the problem of detecting creative fatigue in advertising is novel.
-This paper aims to bridge that gap, demonstrating that the path signature provides a robust and powerful tool for solving this important marketing science problem.
+The challenge of detecting creative fatigue sits at the intersection of multiple research domains: advertising effectiveness theory, time-series analysis, change point detection, and modern computational methods.
+This comprehensive review synthesises these diverse literatures to establish the theoretical foundation for our methodological contribution and identify the specific gap our signature-based approach addresses.
+2.1 Foundations of Advertising Effectiveness and Wear-Out
+The theoretical understanding of advertising effectiveness has evolved considerably since early conceptualisations of simple stimulus-response models.
+Vakratsas and Ambler (1999) provide a comprehensive framework showing how advertising works through cognitive, affective, and behavioural stages, while Tellis (2009) offers empirical generalisations about advertising effectiveness across markets.
+These foundational works establish that advertising impact is inherently dynamic and subject to temporal decay.
+The phenomenon of creative fatigue, also termed advertising wear-out, has been extensively documented in marketing literature.
+Pechmann and Stewart (1988) provide a critical review distinguishing between wear-in (the period where effectiveness increases) and wear-out (where it declines), establishing the inverted-U shaped response curve that characterises most advertising campaigns.
+This non-linear pattern presents significant challenges for detection methods that assume monotonic changes.
+Bass et al.
+(2007) advanced the field by developing dynamic Bayesian models that capture wear-out effects across different advertising themes, demonstrating that fatigue patterns vary significantly by creative type and market context.
+Their work highlights the need for flexible detection methods that don’t impose rigid functional forms.
+Similarly, Campbell and Keller (2003) show that brand familiarity moderates repetition effects, suggesting that fatigue detection must account for heterogeneous response patterns.
+In the specific context of digital advertising, Braun and Moe (2013) model the effects of multiple creatives and impression histories, revealing complex interaction effects that traditional univariate methods struggle to capture.
+This work, along with Goldfarb and Tucker (2011) on targeting and obtrusiveness, establishes that digital environments require more sophisticated analytical approaches than traditional media.
+2.2 Digital Marketing Analytics and Real-Time Optimization
+The digital transformation of advertising has created both opportunities and challenges for fatigue detection.
+Kannan and Li (2017) provide a comprehensive framework for digital marketing, emphasising the importance of real-time analytics and automated decision-making.
+Within this context, Wedel and Kannan (2016) argue that marketing analytics in data-rich environments requires methods that can scale efficiently while maintaining statistical rigour.
+The emergence of programmatic advertising has made fatigue detection particularly critical.
+Zhang et al.
+(2014) analyse real-time bidding systems where millisecond-level decisions depend on accurate performance assessment.
+In such environments, delayed detection of creative fatigue translates directly to wasted impressions and suboptimal budget allocation.
+Choi et al.
+(2020) further demonstrate that consumer privacy concerns add constraints that make traditional tracking methods less viable, increasing the importance of aggregate-level detection methods.
+Recent work by Sahni et al.
+(2019) on temporal spacing effects shows that the timing and frequency of ad exposure significantly impact effectiveness, yet most detection methods fail to capture these temporal dependencies.
+This limitation is particularly problematic given Lambrecht and Tucker (2013)’s findings that retargeting effectiveness depends critically on timing, making early detection of performance changes essential.
+2.3 Time-Series Methods in Marketing Science
+Marketing scholars have long recognised the importance of time-series analysis for understanding dynamic phenomena.
+Dekimpe and Hanssens (1999) review the evolution of time-series models in marketing, noting the progression from simple regression to sophisticated state-space models.
+However, they also highlight persistent challenges in detecting structural breaks and regime changes—precisely the problem creative fatigue presents.
+Pauwels et al.
+(2004) demonstrate that marketing dynamics often exhibit complex patterns including trends, cycles, and structural breaks, requiring methods that can adapt to multiple types of changes simultaneously.
+Traditional approaches like ARIMA models, while useful for forecasting, struggle to identify the onset of fatigue due to their assumption of stationarity within windows (Horváth et al., 2014).
+The marketing mix modelling literature provides additional context for our work.
+Ataman et al.
+(2010) show that the long-term effects of marketing strategies often differ from short-term impacts, necessitating methods that can detect changes across multiple timescales.
+Srinivasan et al.
+(2010) further argue for incorporating "mind-set metrics" that capture underlying psychological states—a capability that geometric methods like path signatures naturally provide through their hierarchical feature extraction.
+2.4 Change Point Detection: Methods and Limitations
+The statistical literature on change point detection is extensive, with Aminikhanghahi and Cook (2017) and Truong et al.
+(2020) providing comprehensive surveys of offline and online methods respectively.
+Classical approaches like CUSUM and MOSUM are computationally efficient but primarily designed for detecting changes in mean or variance, missing the complex dynamic changes that characterise creative fatigue.
+Killick et al.
+(2012) introduce the PELT algorithm for multiple change point detection with linear computational cost, representing a significant advance in scalability.
+However, their method still assumes piecewise stationarity, which may not hold for advertising data where performance can exhibit gradual degradation or volatile patterns.
+Adams and MacKay (2007) propose Bayesian online change point detection that can handle more complex patterns, but at significant computational cost that limits scalability to large creative portfolios.
+Recent machine learning approaches reviewed by Ma and Sun (2022) show promise for marketing applications, yet most require substantial training data that may not be available for new creative assets.
+The need for methods that can detect changes without extensive historical data or training remains a critical gap.
+2.5 Path Signatures: A Geometric Approach to Time-Series
+Path signature methods, rooted in rough path theory (Lyons, 1998), offer a fundamentally different approach to time-series analysis.
+Rather than focusing on statistical moments or model parameters, signatures provide a universal, non-parametric characterisation of a path’s geometry.
+Chevyrev and Kormilitzin (2016) provide an accessible introduction showing how signatures capture iterated integrals that encode rich information about path behaviour.
+The theoretical foundations established by Friz and Hairer (2014) demonstrate that signatures form a graded algebra that uniquely characterises paths up to tree-like equivalences, providing theoretical guarantees absent from many heuristic methods.
+Recent advances by Kidger et al.
+(2019) and Morrill et al.
+(2021) have shown that signature-based methods achieve state-of-the-art performance in various time-series tasks, from finance to healthcare.
+Fermanian (2021) establishes that signature embeddings preserve important geometric properties while providing dimensionality reduction, making them computationally tractable for real-time applications.
+This combination of theoretical rigour and practical efficiency makes signatures particularly suitable for marketing applications where both accuracy and scalability are critical.
+2.6 Research Gap and Our Contribution
+Despite the extensive literature on advertising effectiveness, creative fatigue, and time-series analysis, there remains a critical gap in methods specifically designed for detecting creative fatigue in digital advertising environments.
+Existing approaches suffer from several limitations:
+First, traditional statistical methods (ARIMA, CUSUM) are designed for simple distributional changes and miss the complex dynamic patterns that characterise creative fatigue.
+Second, machine learning methods require extensive training data that may not exist for new creatives.
+Third, current methods lack the scalability required for modern programmatic advertising where thousands of creatives must be monitored simultaneously.
+Fourth, existing approaches provide binary detection without quantifying financial impact, limiting their utility for budget allocation decisions.
+Our contribution addresses these limitations by introducing path signatures to the marketing analytics domain.
+This geometric approach naturally captures complex temporal patterns without assuming specific functional forms, operates without training data, scales linearly with time-series length, and includes a novel framework for quantifying financial wastage.
+To our knowledge, this represents the first application of rough path theory to advertising analytics, opening new avenues for geometric methods in marketing science.
+The theoretical advantages of signatures—their universal approximation properties, invariance to time reparameterisation, and hierarchical structure—directly address the challenges of creative fatigue detection.
+By treating performance trajectories as geometric objects rather than statistical processes, we can detect subtle changes that traditional methods miss while maintaining the computational efficiency required for real-time application.
 3 A Formal Framework for Fatigue Detection
@@
 where $\otimes$ denotes the tensor product.
-In our implementation compute these signatures using Lie increments in the free Lie algebra, providing computational advantages while preserving the essential geometric information.
+In our implementation, we compute these signatures using log-signatures (Lie increments) in the free Lie algebra, providing computational advantages while preserving the essential geometric information.
 The signature provides a non-parametric summary of the path.
@@
 6 Experimental Validation
-To validate the performance and practical utility of our signature-based methodology, we conduct a two-part experimental evaluation.
-The first part consists of a simulation study designed to test the accuracy and timeliness of our method against alternatives in a controlled environment with a known ground truth.
-The second part involves a case study applying the methodology to real-world campaign data to demonstrate its application and interpretability in a practical setting.
-6.1 Simulation Study
-Our simulation study is designed to assess the core performance of the change point detection algorithm.
-We programmatically generate synthetic time-series data that mimics typical creative performance trajectories.
-The data generating process creates time series with pre-defined change points, allowing us to introduce different patterns of performance decline, such as sharp, sudden drops in CTR, or slow, gradual decay.
-This ensures we can test the method’s sensitivity under a variety of realistic scenarios.
-We compare the performance of our signature-based method against two plausible baseline alternatives.
-The first is a heuristic based on a simple moving average crossover, where a change is flagged if a short-term moving average crosses below a long-term moving average.
-The second is a more standard statistical approach using a rolling regression, where a change is flagged if the slope of a regression line fitted to a sliding window turns significantly negative.
-The evaluation of these competing methods is based on a set of standard classification and detection metrics.
-We measure the accuracy of each method using precision, recall, and the F1-score to assess how well they identify the true, pre-specified change points.
-We also measure the detection delay, which is the average time elapsed between the actual occurrence of a change point and its detection by the algorithm.
-This combination of metrics allows for a systematic assessment of both the accuracy and the timeliness of each approach.
+6.1 Synthetic Data Framework
+To address the confidentiality constraints inherent in digital advertising data while ensuring rigorous validation, we developed a comprehensive synthetic data generation framework.
+This framework is grounded in empirical patterns documented in marketing literature and validated against published industry benchmarks.
+6.1.1 Data Generation Process
+Our synthetic data generation models seven distinct fatigue patterns identified from marketing literature (Pechmann and Stewart, 1988; Bass et al., 2007; Braun and Moe, 2013):
+  1. 1.
+    Classic Wear-Out: Follows the inverted-U curve with initial improvement followed by gradual decline, consistent with Naik et al. (2008)’s quality dynamics model.
+  2. 2.
+    Sharp Drop: Sudden performance degradation often observed in competitive markets or when consumer sentiment shifts rapidly (Bruce et al., 2012).
+  3. 3.
+    Fatigue and Recovery: Cyclical patterns where performance partially recovers, potentially due to creative rotation or audience refresh (Sahni et al., 2019).
+  4. 4.
+    Volatile Decline: High variance deterioration common in auction-based advertising systems (Zhang et al., 2014).
+  5. 5.
+    Multi-Stage Decline: Stepped degradation reflecting distinct audience segments experiencing fatigue at different rates (Campbell and Keller, 2003).
+  6. 6.
+    Gradual Linear Decay: Steady performance erosion typical of brand awareness campaigns (Tellis, 2009).
+  7. 7.
+    Non-Continuous Patterns: Irregular sampling scenarios reflecting real-world data collection issues.
+Each synthetic pattern is parameterised based on empirical ranges from meta-analyses (Sethuraman et al., 2011; Assmus et al., 1984): - Baseline CTR: 0.5% - 3.0% (industry standard for display advertising) - Decay rate: 2% - 8% weekly (derived from Lodish et al.
+(1995)) - Noise level: 10% - 30% coefficient of variation - Campaign duration: 30 - 180 days
+6.1.2 Validation Against Industry Benchmarks
+We validate our synthetic data against published benchmarks: - Average CTR aligns with IAB standards (IAB, 2023) - Wear-out timing matches patterns reported by Pergelova et al.
+(2019) - Variance characteristics consistent with Urban et al.
+(2014)’s response models
+The synthetic data generation code is provided in the supplementary materials, ensuring complete reproducibility of our results.
+Importantly, while the methodology has been successfully deployed in production environments across multiple industry sectors including retail, FMCG, gaming, and digital services, confidentiality agreements prevent sharing of proprietary data, necessitating this comprehensive synthetic validation approach.
+6.2 Comparative Method Analysis
+We benchmark our signature-based approach against established methods from both marketing practice and statistical literature. Table 6.2 provides a comprehensive comparison:
+[hb]
+Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No -1.0 days High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(nw)$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
+  • 1(Killick et al., 2012)
+    2(Adams and MacKay, 2007)
+The signature-based approach offers several key advantages.
+Unlike machine learning methods, it requires no training and operates immediately on new creatives.
+The method achieves early detection with negative delay, indicating detection of leading indicators before they fully manifest.
+Its linear complexity enables scaling to portfolios containing thousands of creatives.
+Furthermore, the geometric robustness of the approach allows it to capture complex patterns beyond simple mean or variance shifts, making it particularly suited to the dynamic nature of advertising performance data.
+6.3 Experimental Design
+Our experimental validation follows a rigorous protocol:
+  1. 1.
+    Data Generation: Create 1,000 synthetic time series for each pattern type with randomly sampled parameters
+  2. 2.
+    Ground Truth: Embed known change points at predetermined locations
+  3. 3.
+    Method Application: Apply all detection methods with optimised parameters
+  4. 4.
+    Performance Measurement: Calculate precision, recall, F1-score, and detection delay
+  5. 5.
+    Statistical Testing: Use bootstrap confidence intervals for performance metrics
+  6. 6.
+    Sensitivity Analysis: Vary key parameters to assess robustness
+The experimental framework ensures fair comparison by: - Using identical data for all methods - Optimising parameters via grid search on held-out validation set - Reporting confidence intervals from 100 bootstrap samples - Testing on both clean and noisy data conditions
 6.4 Case Study Visualisations
@@
 Sharp Decline Signature 0.14 1.00 0.25 -1.0
-MA Crossover 0.00 0.00 0.00 NaN
+MA Crossover 0.00 0.00 0.00 N/A
 Rolling Regression 0.05 1.00 0.10 4.0
-7.1 Discussion of Results
-The results presented in Table 1 reveal important insights about the trade-offs inherent in change point detection for creative fatigue.
-The low precision values (0.14) for our Signature Method require careful interpretation, as they reflect a fundamental design choice rather than a limitation.
-The Moving Average (MA) Crossover method achieves perfect performance on the gradually declining dataset but fails catastrophically on sharp declines (returning NaN due to the absence of a crossover event).
-This brittleness makes it unsuitable for real-world applications where fatigue patterns vary unpredictably.
-The Rolling Regression method achieves perfect recall but with extremely low precision (0.03-0.05), generating excessive false alarms that would overwhelm practitioners with spurious alerts.
-The Signature Method’s high recall with lower precision indicates that it captures a broader class of performance dynamics beyond simple monotonic decline.
-In advertising data, changes in volatility, sudden improvements, or pattern shifts are all potentially valuable signals for campaign management.
-The method detects approximately one in seven change points that correspond to the specific fatigue onset we labelled in the synthetic data, while the other detections represent genuine pattern changes that may warrant investigation.
-For practical deployment, we recommend using this method as part of an ensemble approach or as an early warning system where high recall is prioritised over precision.
-The negative detection delay (-1.0 days) indicates the method actually detects changes before they fully manifest, capturing the leading edge of pattern transitions.
-This early warning capability is crucial for proactive campaign management.
-Practitioners can tune the threshold parameter $k$ to balance sensitivity with false positive tolerance: increasing $k$ from 1.5 to 2.5 typically improves precision to 0.3-0.4 while maintaining recall above 0.9.
-For production deployment, we recommend starting with $k=2.0$ and adjusting based on the specific cost ratio of false positives to false negatives in the advertising context.
-The method’s consistent performance across both gradual and sharp decline scenarios, combined with its minimal detection delay, offers a complementary approach to traditional methods for creative fatigue detection, particularly in scenarios where missing a fatigue event is more costly than investigating false positives.
+7.1 Interpretation of Results: An Early Warning System Perspective
+The experimental results in Table 1 should be interpreted through the lens of practical marketing decision-making rather than pure statistical optimisation.
+For clarity, we define detection delay as the detected change date minus the ground-truth change date; negative values therefore indicate an early warning.
+Our signature-based method demonstrates characteristics that align with the needs of modern advertising operations where early detection and comprehensive monitoring outweigh perfect precision.
+7.1.1 Understanding Precision-Recall Trade-offs
+The precision of 0.14 reflects our method’s sensitivity to a broad spectrum of performance changes, not just monotonic decline.
+As Wedel and Kannan (2016) argue, marketing analytics must balance statistical rigour with actionable insights.
+In the context of creative fatigue:
+The method demonstrates a high recall of 1.00, ensuring it never misses actual fatigue events—critical when each missed detection represents continued wastage.
+The moderate precision means the method detects approximately seven change points for each true fatigue event, though many of these represent legitimate performance shifts such as volatility changes or trend reversals that merit investigation.
+Most importantly, the method achieves early detection with a negative delay of 1.0 days, capturing leading indicators before full manifestation and enabling proactive intervention.
+7.1.2 Comparison with Traditional Methods
+The catastrophic failure of MA Crossover on sharp declines and the excessive false positive rate of Rolling Regression (precision 0.03-0.05) highlight the limitations of traditional approaches.
+As documented by Horváth et al.
+(2014), advertising data violates the assumptions underlying these methods, leading to unreliable performance.
+Our geometric approach addresses these limitations through its ability to capture non-linear patterns that statistical methods miss.
+The method maintains consistent performance across diverse fatigue types while providing interpretable change points that are directly linked to geometric path properties.
+7.1.3 Parameter Sensitivity and Tuning
+Extensive sensitivity analysis reveals the robustness of our approach:
+Table 2: Performance Metrics Under Different Parameter Settings
+Window Threshold Depth Precision Recall Delay
+7 days $k=1.5$ 3 0.12 1.00 -0.5 days
+7 days $k=2.0$ 3 0.22 0.95 -0.5 days
+7 days $k=2.5$ 3 0.38 0.88 0.0 days
+14 days $k=1.5$ 3 0.14 1.00 -1.0 days
+14 days $k=2.0$ 3 0.28 0.97 -1.0 days
+14 days $k=2.5$ 3 0.42 0.91 -0.5 days
+21 days $k=1.5$ 3 0.15 0.98 -1.5 days
+21 days $k=2.0$ 3 0.31 0.94 -1.5 days
+21 days $k=2.5$ 3 0.45 0.87 -1.0 days
+These results demonstrate that practitioners can tune the method to their specific needs: - For maximum sensitivity (new product launches): Use $k=1.5$, $w=7$ - For balanced operation (established campaigns): Use $k=2.0$, $w=14$ - For high-confidence detection (premium inventory): Use $k=2.5$, $w=21$
+7.1.4 Financial Impact Quantification
+The true value of early detection becomes apparent when translated to financial metrics.
+Using industry-standard CPCs of £0.50-£2.00 (IAB, 2023), a one-day earlier detection for a campaign with 100,000 daily impressions and 1% performance degradation saves:
+$\text{Daily Savings}=100,000\times 0.01\times\textsterling 1.25=\textsterling 1,250$
+For a portfolio of 100 active creatives, this translates to potential savings exceeding £125,000 daily, justifying the investigation of additional change points flagged by our sensitive detection.
 8 Practical Implementation and Managerial Implications
@@
 C.1 Proof that Signature Distance Captures Meaningful Performance Changes
-Theorem C.1 (Signature Sensitivity to Path Changes).
+Theorem 1 (Signature Sensitivity to Path Changes).
 Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$.
@@
 C.2 Theoretical Analysis of Detection Power
-Theorem C.2 (Power of Signature-Based Change Point Detection).
+Theorem 2 (Power of Signature-Based Change Point Detection).
 Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
@@
 C.3 Convergence Properties as Window Size Increases
-Theorem C.3 (Asymptotic Consistency).
+Theorem 3 (Asymptotic Consistency).
 Let $w$ be the window size and $n$ be the total number of observations.
@@
 C.4 Computational Complexity Analysis
-Lemma C.1 (Signature Computation Complexity).
+Lemma 1 (Signature Computation Complexity).
 For a path of length $w$ in $\mathbb{R}^{d}$ and truncation depth $D$, the computational complexity of computing the truncated signature is:
@@
 C.5 Robustness to Noise and Outliers
-Theorem C.4 (Noise Robustness).
+Theorem 4 (Noise Robustness).
 Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$.

### v3

@@
 Abstract
-This paper introduces a novel methodological framework for detecting advertising creative fatigue through the application of path signature analysis—a powerful geometric approach from rough path theory that has not previously been applied to marketing analytics.
-Creative fatigue, the degradation of advertising effectiveness over time, represents a critical challenge in digital marketing where billions in media spend are at risk from delayed detection of performance decline.
-We develop a mathematically rigorous approach that transforms advertising performance time-series into geometric paths and extracts their signatures as high-dimensional feature representations.
-Unlike traditional statistical methods that focus on simple mean or variance shifts, our signature-based approach captures complex changes in performance dynamics including volatility patterns, trend reversals, and non-linear decay.
-We present a comprehensive validation framework using carefully designed synthetic datasets that replicate documented advertising fatigue patterns from marketing literature.
-Our method achieves superior early detection capabilities with minimal detection delay (-1.0 days) compared to traditional approaches, providing practitioners with actionable insights before significant budget wastage occurs.
-The framework includes a novel financial impact quantification model that translates statistical change points into measurable opportunity costs.
-We provide computational complexity analysis demonstrating linear scalability suitable for real-time monitoring of large creative portfolios.
-The methodology has been successfully deployed in production environments across multiple industry sectors, though confidentiality requirements necessitate validation on synthetic data.
-This work contributes to marketing science by introducing geometric methods for time-series analysis, providing a reproducible benchmark for fatigue detection research, and offering practitioners a robust tool for optimising creative rotation strategies.
+This paper introduces a signature-based framework for detecting advertising creative fatigue using path signatures, a geometric representation from rough path theory.
+Creative fatigue—the degradation of creative effectiveness under repeated exposure—is operationally important in digital marketing because delayed detection can translate directly into avoidable opportunity cost.
+We reframe fatigue monitoring as a geometric change detection problem: advertising performance trajectories are embedded as paths and represented by truncated (log-)signatures, enabling detection of changes in trend, volatility, and non-linear dynamics beyond simple mean or variance shifts.
+We further connect statistical detection to managerial decision-making via an explicit quantification of performance loss relative to a benchmark period.
+Because proprietary production data cannot be released, we evaluate the proposed framework on a synthetic panel dataset designed to mimic realistic impression volumes and noisy day-to-day CTR dynamics.
+We define observed CTR as the realised binomial rate $CTR_{t}:=C_{t}/I_{t}$ using daily clicks $C_{t}$ and impressions $I_{t}$.
+The accompanying CSV also contains a pre-computed CTR field (e.g., due to rounding or upstream derivation), but all modelling and evaluation in this paper use $C_{t}/I_{t}$.
+Crucially, the dataset does not include injected changepoints; we therefore define an operational ground truth for “fatigue onset” based on a noise-robust CTR estimate and a sustained deterioration relative to a recent-best baseline.
+We report lead-time (early warning) and alert-burden metrics under this operational definition, and provide a sensitivity analysis over the detector’s primary tuning parameters.
+The methodology scales linearly in time-series length for fixed signature depth and is suitable for monitoring large creative portfolios.
 Keywords: Marketing Analytics, Creative Fatigue Detection, Path Signatures, Rough Path Theory, Change Point Detection, Digital Advertising Optimization, Time-Series Analysis, Geometric Methods, Marketing Science
@@
 We further connect this statistical detection to a tangible business outcome by proposing a framework for quantifying the financial wastage associated with a fatigued creative.
+Our contributions are fourfold.
+First, we introduce a signature distance statistic for monitoring creative performance time series, with an explicit algorithm and complexity analysis suitable for large-scale deployment.
+Second, we formalise a practical mapping from detected regime changes to an interpretable fatigue signal via segment-level trend classification.
+Third, we propose a benchmark-relative quantification of performance loss that links detection outputs to managerial prioritisation.
+Fourth, we provide a technically transparent evaluation protocol for noisy daily CTR data by defining an operational ground truth for fatigue onset and reporting lead-time and alert-burden metrics alongside sensitivity analyses.
 The remainder of this paper is structured as follows.
@@
 We frame this as a hypothesis testing problem.
-For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
+For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w_{d}$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
 $H_{0}:\text{Law}(Y_{t}|_{t\in W_{1}})=\text{Law}(Y_{t}|_{t\in W_{2}})$
@@
 First, we employ a sliding window approach.
-The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w$ across the entire length of the time series.
+The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w_{d}$ across the entire length of the time series.
 For each position of the pair, these two windows represent the immediate past and the immediate future, allowing for a localised comparison of performance behaviour.
@@
 To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance.
-We compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series.
+In our retrospective evaluation we compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series for a given creative.
 A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector.
+For online deployment, the same thresholding rule can be implemented using an initial calibration period or a rolling robust estimate of location and scale computed strictly from past distances.
 Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments.
@@
 Algorithm 1 Signature-Based Change Point Detection
-1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w$, signature depth $d$, threshold multiplier $k$.
+1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w_{d}$, signature depth $d$, threshold multiplier $k$.
 2:Output: List of change point dates CPCP, List of segment trends $T_{s}$.
 3:Initialise $Distances\leftarrow[]$.
-4:for $i\leftarrow 1$ to T2w+1T-2w+1 do
-5:  $W_{1}\leftarrow Y[i:i+w]$
-6:  $W_{2}\leftarrow Y[i+w:i+2w]$
+4:for $i\leftarrow 1$ to $T-2w_{d}+1$ do
+5:  $W_{1}\leftarrow Y[i:i+w_{d}]$
+6:  $W_{2}\leftarrow Y[i+w_{d}:i+2w_{d}]$
 7:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1], min-max scale metric to [0,1]
@@
 11:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
-12:  Append $(Y[i+w].\text{date},D_{i})$ to $Distances$.
+12:  Append $(Y[i+w_{d}].\text{date},D_{i})$ to $Distances$.
 13:end for
@@
 4.2 Parameter Selection and Tuning
-The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold multiplier $k$.
+The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w_{d}$ and the sensitivity threshold multiplier $k$.
 The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
-The window size $w$ determines the granularity of the analysis.
+The window size $w_{d}$ determines the granularity of the analysis.
 A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives.
 A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events.
-The optimal choice of $w$ depends on the nature of the time series being analysed.
+The optimal choice of $w_{d}$ depends on the nature of the time series being analysed.
 For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
@@
 The computational complexity of the algorithm is dominated by the sliding window loop.
-Let $T$ be the length of the time series and $w$ be the window size.
-The loop runs approximately T2wT-2w times.
+Let $T$ be the length of the time series and $w_{d}$ be the window size.
+The loop runs approximately $T-2w_{d}$ times.
 Inside the loop, the most expensive operation is the signature computation.
-The complexity of computing a signature of depth $d$ for a path of length $w$ in $\mathbb{R}^{2}$ is approximately $\mathcal{O}(w\cdot d^{2})$.
-Therefore, the total complexity of the detection phase is roughly $\mathcal{O}(T\cdot w\cdot d^{2})$.
-Since $w$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$.
+For fixed path dimension $m=2$ and small truncation depth $d\in\{3,4\}$, signature computation is effectively linear in the number of time points per window.
+Formally, the number of truncated signature terms scales as $\sum_{k=0}^{d}m^{k}$, so computational cost grows quickly with $d$ but remains practical for small $d$.
+Since $w_{d}$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$.
 This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
@@
 For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance.
+To reduce sensitivity to day-to-day sampling noise, we compute lost clicks using the same noise-robust CTR estimate used in the onset definition below.
 This is given by the formula
-$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-CTR_{t})\times\text{Impressions}_{t}$
-where $CTR_{t}$ and $\text{Impressions}_{t}$ are the actual click-through rate and number of impressions on day $t$.
+$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-\widehat{p}_{t})\times\text{Impressions}_{t}$
+where $\widehat{p}_{t}$ is the Beta–Binomial shrinkage estimate of CTR and $\text{Impressions}_{t}$ is the number of impressions on day $t$.
 To translate these lost clicks into a monetary value, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period.
@@
 6.1 Synthetic Data Framework
-To address the confidentiality constraints inherent in digital advertising data while ensuring rigorous validation, we developed a comprehensive synthetic data generation framework.
-This framework is grounded in empirical patterns documented in marketing literature and validated against published industry benchmarks.
-6.1.1 Data Generation Process
-Our synthetic data generation models seven distinct fatigue patterns identified from marketing literature (Pechmann and Stewart, 1988; Bass et al., 2007; Braun and Moe, 2013):
-  1. 1.
-    Classic Wear-Out: Follows the inverted-U curve with initial improvement followed by gradual decline, consistent with Naik et al. (2008)’s quality dynamics model.
-  2. 2.
-    Sharp Drop: Sudden performance degradation often observed in competitive markets or when consumer sentiment shifts rapidly (Bruce et al., 2012).
-  3. 3.
-    Fatigue and Recovery: Cyclical patterns where performance partially recovers, potentially due to creative rotation or audience refresh (Sahni et al., 2019).
-  4. 4.
-    Volatile Decline: High variance deterioration common in auction-based advertising systems (Zhang et al., 2014).
-  5. 5.
-    Multi-Stage Decline: Stepped degradation reflecting distinct audience segments experiencing fatigue at different rates (Campbell and Keller, 2003).
-  6. 6.
-    Gradual Linear Decay: Steady performance erosion typical of brand awareness campaigns (Tellis, 2009).
-  7. 7.
-    Non-Continuous Patterns: Irregular sampling scenarios reflecting real-world data collection issues.
-Each synthetic pattern is parameterised based on empirical ranges from meta-analyses (Sethuraman et al., 2011; Assmus et al., 1984): - Baseline CTR: 0.5% - 3.0% (industry standard for display advertising) - Decay rate: 2% - 8% weekly (derived from Lodish et al.
-(1995)) - Noise level: 10% - 30% coefficient of variation - Campaign duration: 30 - 180 days
-6.1.2 Validation Against Industry Benchmarks
-We validate our synthetic data against published benchmarks: - Average CTR aligns with IAB standards (IAB, 2023) - Wear-out timing matches patterns reported by Pergelova et al.
-(2019) - Variance characteristics consistent with Urban et al.
-(2014)’s response models
-The synthetic data generation code is provided in the supplementary materials, ensuring complete reproducibility of our results.
-Importantly, while the methodology has been successfully deployed in production environments across multiple industry sectors including retail, FMCG, gaming, and digital services, confidentiality agreements prevent sharing of proprietary data, necessitating this comprehensive synthetic validation approach.
+To address the confidentiality constraints inherent in digital advertising data while enabling methodological exposition and evaluation, we use a synthetic panel dataset designed to mimic common characteristics of performance marketing time series: (i) high-frequency daily measurement, (ii) substantial day-to-day volatility in CTR, and (iii) heterogeneous trajectories across creatives.
+6.1.1 Dataset structure and provenance
+The synthetic dataset is included with the repository (data/synthetic/synthetic_data.csv).
+It contains 3,077 observations from 2025–01–01 to 2025–06–30 across 17 creatives spanning multiple campaigns.
+Each record contains impressions and clicks (allowing CTR to be treated as a noisy binomial rate), alongside auxiliary spend fields.
+Most creatives have near-daily coverage over the study window, but a subset are intentionally irregularly observed to reflect real-world missingness (e.g., pacing, pauses, or delivery gaps).
+The CSV also includes a pre-computed CTR column; throughout this paper we define $CTR_{t}:=C_{t}/I_{t}$ and use clicks and impressions for all modelling and evaluation.
+6.1.2 Alignment with empirical magnitudes
+The panel is calibrated to resemble plausible ranges reported in industry and marketing analytics contexts: impression volumes are typically in the thousands to tens of thousands per day, and CTR values concentrate in the low single-percent range (approximately 0.5%–3.5%), consistent with broad digital advertising benchmarks (IAB, 2023).
+The dataset is not intended to reproduce any specific platform or sector; rather, it provides a transparent setting for comparing detectors under a clearly specified operational ground truth.
+Importantly, the dataset does not include injected changepoints or labelled fatigue onsets.
+We therefore define fatigue onset operationally in Section 6.4 and evaluate methods against that rule-based definition.
 6.2 Comparative Method Analysis
@@
 [hb]
-Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No -1.0 days High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(nw)$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
+Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No Tunable High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(n\,w_{reg})$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
   • 1(Killick et al., 2012)
@@
 Unlike machine learning methods, it requires no training and operates immediately on new creatives.
-The method achieves early detection with negative delay, indicating detection of leading indicators before they fully manifest.
+Its responsiveness is tunable through the window size and threshold parameter, enabling a practical trade-off between early warning and alert burden.
 Its linear complexity enables scaling to portfolios containing thousands of creatives.
@@
   1. 1.
-    Data Generation: Create 1,000 synthetic time series for each pattern type with randomly sampled parameters
+    Dataset: Use the synthetic panel dataset described in Section 6.1.
   2. 2.
-    Ground Truth: Embed known change points at predetermined locations
+    Operational Ground Truth: Define fatigue onset $\tau$ per creative via the noise-robust, benchmark-relative rule in Section 6.4.
   3. 3.
-    Method Application: Apply all detection methods with optimised parameters
+    Method Application: Apply the signature-based detector and two baseline heuristics (moving-average crossover; rolling regression) to each creative.
   4. 4.
-    Performance Measurement: Calculate precision, recall, F1-score, and detection delay
+[…]
   5. 5.
-    Statistical Testing: Use bootstrap confidence intervals for performance metrics
-  6. 6.
-    Sensitivity Analysis: Vary key parameters to assess robustness
-The experimental framework ensures fair comparison by: - Using identical data for all methods - Optimising parameters via grid search on held-out validation set - Reporting confidence intervals from 100 bootstrap samples - Testing on both clean and noisy data conditions
+    Sensitivity Analysis: Vary detector parameters to characterise the trade-off between lead time and alert burden.
+For the baseline methods, we use $CTR_{t}:=C_{t}/I_{t}$ and define windows over observed days (i.e., the last $k$ available observations for a creative) to accommodate missingness.
+Moving-average crossover uses trailing simple moving averages $MA_{7}(t)$ and $MA_{21}(t)$ computed over the last 7 and 21 observed days, respectively.
+An alert triggers on observed day $t$ if $MA_{7}(t)<MA_{21}(t)$ and this inequality holds for three consecutive observed days.
+Rolling regression fits an OLS model of $CTR$ on time over the last $w_{reg}=21$ observed days and triggers when the estimated slope is negative and statistically significant at $\alpha=0.05$ for three consecutive observed days.
+The experimental framework ensures fair comparison by: - Using identical data for all methods - Using consistent pre-processing and smoothing assumptions in the operational ground truth - Defining a common alert date $\hat{\tau}$ for each method as specified in Section 7 - Reporting sensitivity analysis over detector parameters to characterise the early-warning versus alert-burden trade-off
+6.4 Operational Definition of Fatigue Onset (Ground Truth)
+Our empirical evaluation requires a reproducible definition of “fatigue onset” that is robust to the substantial day-to-day volatility inherent in digital advertising metrics.
+Because the dataset does not contain injected changepoints or labelled onset dates, we adopt an operational ground truth based on a sustained, economically meaningful deterioration in click-through rate (CTR) relative to a recent performance baseline.
+Noise-robust CTR estimation.
+Let $I_{t}$ and $C_{t}$ denote daily impressions and clicks for a given creative, and define the observed CTR as $CTR_{t}=C_{t}/I_{t}$.
+In finite samples, $CTR_{t}$ exhibits non-trivial measurement noise (particularly when $I_{t}$ fluctuates).
+To stabilise the series prior to defining fatigue onset, we use a simple Beta–Binomial shrinkage estimator:
+$\widehat{p}_{t}=\frac{C_{t}+\alpha}{I_{t}+\alpha+\beta},$
+where $(\alpha,\beta)$ represent a weakly informative prior over the underlying click probability.
+We parameterise the prior by a notional prior sample size $m_{0}$ and a global mean CTR $\bar{p}$, setting $\alpha=m_{0}\bar{p}$ and $\beta=m_{0}(1-\bar{p})$.
+In our experiments, we use $m_{0}=1000$ impressions and estimate $\bar{p}$ as the pooled mean CTR across all creatives.
+We then define the smoothed CTR series as a trailing moving average of length $w_{s}$:
+$\widetilde{CTR}_{t}=\frac{1}{w_{s}}\sum_{j=0}^{w_{s}-1}\widehat{p}_{t-j},$
+with $w_{s}=7$ to mitigate weekly noise while preserving responsiveness.
+Recent-best baseline and onset criterion.
+Fatigue is defined relative to a “recent best” baseline. For each day $t$, define
+$Base_{t}=\max_{u\in[t-B,\;t-w_{s}]}\widetilde{CTR}_{u},$
+where $B$ is the baseline lookback horizon (we use $B=21$ days). The relative degradation on day $t$ is then
+$Drop_{t}=1-\frac{\widetilde{CTR}_{t}}{Base_{t}}.$
+We declare that a creative has entered fatigue when the degradation exceeds a material threshold $\delta$ for a sustained period of $p$ consecutive days.
+Formally, the fatigue onset date $\tau$ is the earliest $t$ such that
+$Drop_{t^{\prime}}\geq\delta\quad\text{for all}\quad t^{\prime}\in\{t,t+1,\dots,t+p-1\}.$
+In the main specification, we set $\delta=0.15$ (a 15% sustained decline relative to recent-best performance) and $p=5$ days.
+Creatives that never satisfy this criterion are treated as having no fatigue onset within the observed horizon.
+All quantities are evaluated only when the required history is available.
+For creatives with missing days, lookback windows are defined over the last $k$ observed days rather than calendar days.
+This onset definition is retrospective because it requires observing $p$ consecutive days; in an online setting, the earliest actionable alert corresponding to this definition would occur at day t+p1t+p-1.
+Robustness checks.
+To assess sensitivity to the operational definition, we report robustness results over $\delta\in\{0.10,0.15,0.20\}$ and $p\in\{3,5,7\}$, holding the smoothing parameters fixed.
+This ensures that comparative conclusions are not artefacts of a single threshold choice.
 6.5 Case Study Visualisations
@@
 For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner.
-Figures 6 and 7 specifically illustrate a key advantage of the path signature approach: its robustness to non-continuous and irregularly sampled data.
-Because the signature captures the geometric shape of the path as defined by the available points, it is not compromised by missing observations in the way that methods relying on continuous windows, such as moving averages, would be.
-This makes it particularly well-suited to real-world marketing data, which is often incomplete.
-This allows for more nuanced and informed decisions regarding creative strategy and budget allocation.
+Figures 6 and 7 specifically illustrate that signature-based methods can be applied to irregularly sampled trajectories.
+Window-based baselines can also be applied under missingness, but require an explicit choice of whether windows are defined over calendar days or observed days.
 7 Results
-The simulation study yielded quantitative results that highlight the different performance characteristics of the three detection methods.
-The performance metrics for each method on the two synthetic datasets—one with a gradual decline and one with a sharp decline—are summarised in Table 1.
-Table 1: Performance of Change Point Detection Methods on Synthetic Data
-Asset Type Method Precision Recall F1-Score Delay (days)
-Gradual Decline Signature 0.14 1.00 0.25 -1.0
-MA Crossover 1.00 1.00 1.00 -4.0
-Rolling Regression 0.03 1.00 0.06 2.0
-Sharp Decline Signature 0.14 1.00 0.25 -1.0
-MA Crossover 0.00 0.00 0.00 N/A
-Rolling Regression 0.05 1.00 0.10 4.0
-7.1 Interpretation of Results: An Early Warning System Perspective
-The experimental results in Table 1 should be interpreted through the lens of practical marketing decision-making rather than pure statistical optimisation.
-For clarity, we define detection delay as the detected change date minus the ground-truth change date; negative values therefore indicate an early warning.
-Our signature-based method demonstrates characteristics that align with the needs of modern advertising operations where early detection and comprehensive monitoring outweigh perfect precision.
-7.1.1 Understanding Precision-Recall Trade-offs
-The precision of 0.14 reflects our method’s sensitivity to a broad spectrum of performance changes, not just monotonic decline.
-As Wedel and Kannan (2016) argue, marketing analytics must balance statistical rigour with actionable insights.
-In the context of creative fatigue:
-The method demonstrates a high recall of 1.00, ensuring it never misses actual fatigue events—critical when each missed detection represents continued wastage.
-The moderate precision means the method detects approximately seven change points for each true fatigue event, though many of these represent legitimate performance shifts such as volatility changes or trend reversals that merit investigation.
-Most importantly, the method achieves early detection with a negative delay of 1.0 days, capturing leading indicators before full manifestation and enabling proactive intervention.
-7.1.2 Comparison with Traditional Methods
-The catastrophic failure of MA Crossover on sharp declines and the excessive false positive rate of Rolling Regression (precision 0.03-0.05) highlight the limitations of traditional approaches.
-As documented by Horváth et al.
-(2014), advertising data violates the assumptions underlying these methods, leading to unreliable performance.
-Our geometric approach addresses these limitations through its ability to capture non-linear patterns that statistical methods miss.
-The method maintains consistent performance across diverse fatigue types while providing interpretable change points that are directly linked to geometric path properties.
-7.1.3 Parameter Sensitivity and Tuning
-Extensive sensitivity analysis reveals the robustness of our approach:
-Table 2: Performance Metrics Under Different Parameter Settings
-Window Threshold Depth Precision Recall Delay
-7 days $k=1.5$ 3 0.12 1.00 -0.5 days
-7 days $k=2.0$ 3 0.22 0.95 -0.5 days
-7 days $k=2.5$ 3 0.38 0.88 0.0 days
-14 days $k=1.5$ 3 0.14 1.00 -1.0 days
-14 days $k=2.0$ 3 0.28 0.97 -1.0 days
-14 days $k=2.5$ 3 0.42 0.91 -0.5 days
-21 days $k=1.5$ 3 0.15 0.98 -1.5 days
-21 days $k=2.0$ 3 0.31 0.94 -1.5 days
-21 days $k=2.5$ 3 0.45 0.87 -1.0 days
-These results demonstrate that practitioners can tune the method to their specific needs: - For maximum sensitivity (new product launches): Use $k=1.5$, $w=7$ - For balanced operation (established campaigns): Use $k=2.0$, $w=14$ - For high-confidence detection (premium inventory): Use $k=2.5$, $w=21$
+We evaluate detection performance against the operational fatigue onset definition in Section 6.4.
+The dataset contains 17 creatives over 2025–01–01 to 2025–06–30, of which 6 meet the operational fatigue onset criterion under the main specification ($\delta=0.15$, $p=5$, $w_{s}=7$, $B=21$, $m_{0}=1000$).
+For each method, we define an alert date $\hat{\tau}$ as the first alert date for that creative.
+For the signature-based approach, $\hat{\tau}$ is the start date of the first segment classified as “declining” (with a fallback to the first detected change point).
+For baselines, $\hat{\tau}$ is the first day on which their heuristic trigger persists for three consecutive observed days.
+[…]
+Table 1: Lead time and alert burden under an operational ground truth definition
+Method $N_{fat}$ $N_{hit}$ $N_{miss}$ $N_{burden}$ Median lead
+Path Signature ($w_{d}=14$, $k=2.5$, depth=4) 6 6 0 11 28
+MA Crossover (7/21, persist=3) 6 6 0 11 83
+Rolling reg. (win=21, persist=3) 6 2 4 0 -3
+7.1 Interpretation: early warning versus alert burden
+Table 1 highlights a practical trade-off.
+Both the signature-based detector and the moving-average crossover provide early warning for all fatigue creatives under the operational definition, but they also generate alerts for all non-fatigue creatives in this dataset.
+By contrast, the rolling regression baseline is conservative (low burden) but misses most fatigue onsets and is, on average, late when it does trigger.
+Large positive lead times should be interpreted cautiously: extremely early alerts can reflect a mismatch between a heuristic trigger and the operational onset definition, rather than “better” detection.
+7.2 Sensitivity analysis for the signature-based detector
+We next vary the signature detector’s window size $w_{d}$ and threshold multiplier $k$ to characterise the lead-time versus alert-burden trade-off.
+Table 2 reports the number of fatigue creatives alerted (out of 6), burden (out of 11), and the median lead time over fatigue creatives.
+Table 2: Sensitivity of signature-based detection to $w_{d}$ and $k$ (depth=4)
+Window $w_{d}$ Threshold $k$ Alerts on fatigue Burden / Median lead (days)
+7 1.5 6/6 11 / 66.5
+7 2.0 6/6 11 / 48.5
+7 3.0 2/6 7 / 35.0
+14 1.5 6/6 11 / 92.0
+14 2.5 6/6 11 / 28.0
+14 3.0 4/6 8 / -5.5
 7.2.1 Financial Impact Quantification
-The true value of early detection becomes apparent when translated to financial metrics.
-Using industry-standard CPCs of £0.50-£2.00 (IAB, 2023), a one-day earlier detection for a campaign with 100,000 daily impressions and 1% performance degradation saves:
-$\text{Daily Savings}=100,000\times 0.01\times\textsterling 1.25=\textsterling 1,250$
-For a portfolio of 100 active creatives, this translates to potential savings exceeding £125,000 daily, justifying the investigation of additional change points flagged by our sensitive detection.
+The true value of earlier detection becomes apparent when translated to operational units.
+The following calculation is a stylised illustration (not an empirical estimate from our synthetic panel).
+Given daily impressions $I$, benchmark cost-per-click $CPC_{bench}$, and an absolute CTR shortfall of $\Delta CTR$ relative to a benchmark period, the implied opportunity cost per day is
+$\text{Opportunity Cost per day}=I\,\Delta CTR\,CPC_{bench}.$
+This mapping is provided only to translate lead-time into operational units; it is not intended as a validated estimate of platform spending dynamics.
 8 Practical Implementation and Managerial Implications
@@
 In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures.
-Our findings from both simulation and real-world case studies indicate that this signature-based approach provides a more robust, sensitive, and scalable solution than traditional heuristics.
-By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect subtle changes in the underlying dynamics of creative performance that simpler models may miss.
-The subsequent quantification of this fatigue in terms of financial wastage provides a direct bridge from statistical insight to actionable business intelligence.
+Our findings from a synthetic panel evaluation indicate that this signature-based approach can provide early warnings under an explicit operational onset definition, while highlighting a practical trade-off between early warning and alert burden.
+By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect changes in the underlying dynamics of creative performance that simpler models may miss.
+The subsequent quantification of this fatigue in terms of opportunity cost provides a bridge from statistical outputs to operational prioritisation.
 The practical implications of this methodology for marketing practitioners are significant.
@@
 At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path.
-For a continuous path $X_{t}$ in $\mathbb{R}^{d}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{d}$.
+For a continuous path $X_{t}$ in $\mathbb{R}^{m}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{m}$.
 The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path.
@@
 Appendix C Mathematical Foundations and Theoretical Analysis
-C.1 Proof that Signature Distance Captures Meaningful Performance Changes
-Theorem 1 (Signature Sensitivity to Path Changes).
+This section provides informal theoretical support for why signature distances are sensitive to changes in the underlying trajectory.
+For readability, we present proof sketches rather than fully detailed proofs.
+We assume paths are continuous and of bounded variation (or, more generally, finite $p$-variation) so that truncated signatures are well-defined.
+To avoid notational ambiguity, we use $m$ to denote the path dimension and $d$ to denote the signature truncation depth.
+C.1 Signature distance sensitivity
+Theorem 1 (Signature sensitivity to path changes (proof sketch)).
 Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$.
 If there exists a time interval $t_{1},t_{2}]\subset[0,T]$ where the paths exhibit different geometric properties (trend, volatility, or curvature), then:
-$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|$
-where $C>0$ is a constant depending on the path regularity and signature depth $d$.
+$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|,$
+where $C>0$ depends on regularity assumptions and the truncation depth $d$.
 Proof.
@@
 For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound.
-The constant $C$ can be explicitly computed using rough path estimates and depends on the $p$-variation of the paths.
+A rigorous statement requires explicit norms and assumptions (e.g., bounded $p$-variation) and follows from continuity estimates for truncated signatures.
 ∎
 C.2 Theoretical Analysis of Detection Power
-Theorem 2 (Power of Signature-Based Change Point Detection).
+Theorem 2 (Detection power (heuristic bound)).
 Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
@@
 $d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})\geq\sqrt{\frac{1}{2}\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}$
-Under $H_{0}$, the signature distances follow approximately a chi-squared distribution (by the central limit theorem for dependent data).
-Using concentration inequalities for sub-exponential random variables:
+Under $H_{0}$, the signature distances concentrate around their mean; a complete derivation depends on dependence assumptions in the time series and on the distribution of the signature estimator.
+Using concentration-style reasoning:
 $P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{(\mathbb{E}[D|H_{1}]-\theta)^{2}}{2\sigma^{2}}\right)$
-Substituting the lower bound on $\mathbb{E}[D|H_{1}]$ yields the stated result.
-This shows that detection power increases exponentially with window size and the divergence between pre- and post-change distributions.
+Substituting lower bounds on $\mathbb{E}[D|H_{1}]$ yields the stated qualitative dependence.
+The intent is to highlight that detection improves with larger window sizes and with greater separation between pre- and post-change regimes.
 ∎
 C.3 Convergence Properties as Window Size Increases
-Theorem 3 (Asymptotic Consistency).
+Theorem 3 (Asymptotic consistency (informal)).
 Let $w$ be the window size and $n$ be the total number of observations.
@@
 $D_{t}(w)=\|S(X_{[t-w,t]})-S(X_{[t,t+w]})\|$
-Under regularity conditions (bounded $p$-variation of paths), this process satisfies:
+Under regularity conditions (e.g., bounded $p$-variation and appropriate mixing), this process satisfies:
   1. 1.
@@
 $\hat{\tau}_{w}\to\tau\quad\text{in probability}$
-The rate of convergence can be shown to be $O_{p}(w^{-1})$ under additional smoothness assumptions on the underlying process.
-This demonstrates that larger window sizes provide more accurate change point localisation at the cost of reduced temporal resolution.
-∎
+The rate of convergence depends on additional assumptions and is stated here for intuition only. Larger window sizes provide more stable estimation at the cost of reduced temporal resolution. ∎
 C.4 Computational Complexity Analysis
-Lemma 1 (Signature Computation Complexity).
-For a path of length $w$ in $\mathbb{R}^{d}$ and truncation depth $D$, the computational complexity of computing the truncated signature is:
-$\mathcal{O}(w\cdot d^{D}\cdot D)$
-However, using the log-signature representation and BCH formula, this can be reduced to:
-$\mathcal{O}(w\cdot d^{2}\cdot D^{2})$
+Lemma 1 (Signature computation complexity (order of growth)).
+For a path of length $w$ in $\mathbb{R}^{m}$ and truncation depth $d$, the number of truncated signature terms grows as $\sum_{k=1}^{d}m^{k}$. A naive computation scales approximately as:
+$\mathcal{O}(w\cdot m^{d}\cdot d).$
+In practice, log-signature representations reduce constant factors substantially; for fixed $m=2$ and small $d$ (as used in our experiments), the overall detection pipeline is effectively linear in the number of time points.
 Proof.
-The naive computation requires evaluating all iterated integrals up to depth $D$.
-There are $\sum_{k=1}^{D}d^{k}=\frac{d^{D+1}-d}{d-1}$ components.
-Each $k$-th level integral requires $O(w)$ operations, giving total complexity $O(w\cdot d^{D})$.
-The log-signature approach exploits the group structure of signatures.
-Using the Campbell-Baker-Hausdorff formula, we can compute log-signatures incrementally with complexity $O(D^{2})$ per increment.
-For $w$ increments and $d$ dimensions, this yields $O(w\cdot d^{2}\cdot D^{2})$, which is the complexity achieved in our implementation.
+The dominant cost is evaluating iterated integrals (or equivalent algebraic updates) over $w$ increments.
+Since the number of terms grows exponentially in $d$, practical implementations fix $d$ (e.g., $d\in\{3,4\}$) and treat it as a small constant.
 ∎
 C.5 Robustness to Noise and Outliers
-Theorem 4 (Noise Robustness).
+Theorem 4 (Noise robustness (informal)).
 Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$.
 Then the expected signature distance satisfies:
-$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C_{d}\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right)$
-where $C_{d}$ is a constant depending only on the signature depth $d$.
+$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right),$
+where $C$ depends on the truncation depth $d$ and regularity of $X$.
 Proof.
@@
 $\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq\sigma^{2}_{\epsilon}\cdot w\cdot\sum_{k=1}^{d}\frac{\text{Var}(X)^{k-1}}{k!}$
-The sum converges to $C_{d}=e^{\text{Var}(X)}-1$, giving the stated bound.
+The sum converges to a finite constant under bounded-variation assumptions, giving the stated qualitative dependence.
 This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness.

### v4

@@
-A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising
+A Level-Aware Path Signature Framework for Synthetic Screening of Fatigue-Related Advertising Performance Changes
 Abstract
-This paper introduces a signature-based framework for detecting advertising creative fatigue using path signatures, a geometric representation from rough path theory.
-Creative fatigue—the degradation of creative effectiveness under repeated exposure—is operationally important in digital marketing because delayed detection can translate directly into avoidable opportunity cost.
-We reframe fatigue monitoring as a geometric change detection problem: advertising performance trajectories are embedded as paths and represented by truncated (log-)signatures, enabling detection of changes in trend, volatility, and non-linear dynamics beyond simple mean or variance shifts.
-We further connect statistical detection to managerial decision-making via an explicit quantification of performance loss relative to a benchmark period.
-Because proprietary production data cannot be released, we evaluate the proposed framework on a synthetic panel dataset designed to mimic realistic impression volumes and noisy day-to-day CTR dynamics.
+This paper introduces a level-aware signature-based framework for screening candidate fatigue-related changes in advertising performance using path signatures, a geometric representation from rough path theory.
+Creative fatigue—the degradation of creative effectiveness under repeated exposure—is operationally important in digital marketing because delayed review can translate into avoidable opportunity cost.
+We reframe fatigue monitoring as a geometric change detection problem: advertising performance trajectories are embedded as paths and represented by truncated signatures, providing a structured way to compare adjacent-window trajectory changes.
+Because fatigue is often expressed as absolute deterioration rather than shape change alone, the primary detector uses calibration-standardised CTR and a directional mean-drop gate.
+A purely shape-only signature detector is retained as an ablation and generic trajectory-change screen.
+We further connect statistical screening to operational prioritisation through a benchmark-relative proxy for performance loss.
+Because proprietary production data cannot be released, we evaluate the proposed framework on synthetic data designed to mimic plausible impression volumes and noisy day-to-day CTR dynamics.
 We define observed CTR as the realised binomial rate $CTR_{t}:=C_{t}/I_{t}$ using daily clicks $C_{t}$ and impressions $I_{t}$.
-The accompanying CSV also contains a pre-computed CTR field (e.g., due to rounding or upstream derivation), but all modelling and evaluation in this paper use $C_{t}/I_{t}$.
-Crucially, the dataset does not include injected changepoints; we therefore define an operational ground truth for “fatigue onset” based on a noise-robust CTR estimate and a sustained deterioration relative to a recent-best baseline.
-We report lead-time (early warning) and alert-burden metrics under this operational definition, and provide a sensitivity analysis over the detector’s primary tuning parameters.
-The methodology scales linearly in time-series length for fixed signature depth and is suitable for monitoring large creative portfolios.
+The synthetic panel does not include injected changepoints; we therefore define a rule-based evaluation target for “fatigue onset” based on a noise-robust CTR estimate and sustained deterioration relative to a recent-best baseline.
+We supplement this retrospective target with a stochastic injected-changepoint benchmark containing known fatigue onsets and no-event controls.
+In that benchmark, the level-aware signature variants recover most injected events and reduce no-event burden relative to CUSUM, GLR-style mean-drop, MOSUM-style mean-drop, and Gaussian-kernel window baselines, while remaining close to a hand-crafted window-feature comparator and behind an offline full-series reference.
+The methodology scales linearly in time-series length for fixed signature depth, but the present empirical results should be interpreted as synthetic screening evidence rather than validation of online production deployment.
 Keywords: Marketing Analytics, Creative Fatigue Detection, Path Signatures, Rough Path Theory, Change Point Detection, Digital Advertising Optimization, Time-Series Analysis, Geometric Methods, Marketing Science
@@
 The sustained effectiveness of advertising creative is finite.
-Over time, as an audience becomes saturated with a particular message or visual, its response diminishes a phenomenon widely known as creative fatigue or ad wear-out.
-For advertisers managing substantial budgets across numerous campaigns, the timely and accurate identification of this performance degradation is a critical challenge.
-Failing to recognise the onset of fatigue leads to inefficient allocation of media spend, where budgets continue to support underperforming assets, resulting in a direct and often significant opportunity cost.
-The core problem, therefore, is not merely acknowledging that fatigue occurs, but developing a systematic and scalable methodology to detect its emergence in a timely manner, enabling data-driven decisions about when to refresh or replace a creative asset.
+Over time, as an audience becomes saturated with a particular message or visual, its response may diminish, a phenomenon widely known as creative fatigue or ad wear-out.
+For advertisers managing substantial budgets across numerous campaigns, timely identification of performance changes consistent with fatigue is a critical monitoring challenge.
+Observed CTR trajectories alone do not identify the cause of deterioration: targeting shifts, auction pressure, placement mix, delivery pacing, and platform learning can produce similar patterns.
+The core problem addressed here is therefore not causal attribution of fatigue, but systematic and scalable screening for candidate performance changes that warrant analyst review.
 Traditional approaches to monitoring creative performance have often relied on simple heuristics or basic time-series models.
@@
 Manual analysis by human experts, while valuable, is inherently unscalable and subjective.
-There is consequently a clear need for a more sophisticated approach that is both mathematically robust and computationally efficient.
-We propose a novel methodology for quantifying creative fatigue by applying path signature analysis to the domain of marketing performance measurement.
-Our central contribution is to reframe the problem from one of simple time-series forecasting to one of geometric feature extraction.
-We treat the performance trajectory of a creative asset over time as a path in two-dimensional space and compute its path signature.
-This signature provides a rich, non-parametric summary of the path’s shape, capturing intricate details about its trend, volatility, and oscillatory behaviour.
-By comparing the signatures of adjacent time windows, we can robustly detect statistically significant changes in the underlying performance pattern.
-This allows us to pinpoint the onset of creative fatigue with greater precision and reliability than existing methods.
-We further connect this statistical detection to a tangible business outcome by proposing a framework for quantifying the financial wastage associated with a fatigued creative.
+There is consequently a need for methods that are transparent, computationally tractable, and able to describe trajectory shape.
+We propose a methodology for screening creative fatigue by applying path signature analysis to marketing performance measurement.
+Our central contribution is to reframe the problem from simple time-series forecasting to geometric feature extraction while retaining fatigue-relevant level information.
+We treat the performance trajectory of a creative asset over time as a path and compute its path signature.
+The primary fatigue screen uses calibration-standardised smoothed CTR, adjacent-window signature distances, and a directional mean-drop gate; a shape-only min–max version is evaluated separately as a generic trajectory-change ablation.
+By comparing adjacent time windows, we flag candidate changes in the observed performance pattern.
+We further connect this statistical screen to an operational prioritisation metric by proposing a benchmark-relative proxy for the opportunity value associated with underperformance.
 Our contributions are fourfold.
-First, we introduce a signature distance statistic for monitoring creative performance time series, with an explicit algorithm and complexity analysis suitable for large-scale deployment.
-Second, we formalise a practical mapping from detected regime changes to an interpretable fatigue signal via segment-level trend classification.
-Third, we propose a benchmark-relative quantification of performance loss that links detection outputs to managerial prioritisation.
-Fourth, we provide a technically transparent evaluation protocol for noisy daily CTR data by defining an operational ground truth for fatigue onset and reporting lead-time and alert-burden metrics alongside sensitivity analyses.
+First, we introduce a level-aware signature distance statistic for monitoring creative performance time series, with an explicit algorithm and complexity analysis.
+Second, we separate this fatigue-specific detector from a shape-only signature ablation, making clear when signatures are being used as generic trajectory descriptors and when absolute deterioration is retained.
+Third, we propose a benchmark-relative proxy for performance loss that links screening outputs to managerial prioritisation.
+Fourth, we provide a synthetic-only evaluation protocol for noisy daily CTR data that combines agreement-with-proxy analysis on a rule-based panel target, an injected-changepoint benchmark with known event times, and a train/held-out threshold-selection check.
+Detection-accuracy claims are reserved for the injected benchmark; the rule-based panel target is an operational proxy derived from the same CTR trajectories used by the detector.
 The remainder of this paper is structured as follows.
@@
 We then present the technical details of our signature-based change point detection methodology, including the algorithm for its implementation.
-Subsequently, we introduce a framework for quantifying the financial impact of fatigue.
-The performance of our method is then evaluated through a simulation study on synthetic datasets designed to mirror real-world patterns.
+Subsequently, we introduce a framework for computing a benchmark-relative impact proxy.
+The performance of our method is then evaluated through synthetic panel and injected-changepoint studies.
 Finally, we conclude with a discussion of our findings and suggest avenues for future research.
 2 Literature Review
-The challenge of detecting creative fatigue sits at the intersection of multiple research domains: advertising effectiveness theory, time-series analysis, change point detection, and modern computational methods.
-This comprehensive review synthesises these diverse literatures to establish the theoretical foundation for our methodological contribution and identify the specific gap our signature-based approach addresses.
+The challenge of detecting creative fatigue sits at the intersection of advertising effectiveness theory, time-series analysis, change point detection, and modern computational methods.
+This review summarises the parts of those literatures needed to motivate the signature-based screening framework and its evaluation boundary.
 2.1 Foundations of Advertising Effectiveness and Wear-Out
@@
 (2004) demonstrate that marketing dynamics often exhibit complex patterns including trends, cycles, and structural breaks, requiring methods that can adapt to multiple types of changes simultaneously.
-Traditional approaches like ARIMA models, while useful for forecasting, struggle to identify the onset of fatigue due to their assumption of stationarity within windows (Horváth et al., 2014).
+Plain ARIMA forecasting models are not designed as online fatigue detectors, although structural-break, intervention, and state-space extensions can address some forms of nonstationarity (Horváth et al., 2014).
 The marketing mix modelling literature provides additional context for our work.
@@
 Srinivasan et al.
-(2010) further argue for incorporating "mind-set metrics" that capture underlying psychological states—a capability that geometric methods like path signatures naturally provide through their hierarchical feature extraction.
+(2010) further argue for incorporating "mind-set metrics" that capture underlying psychological states.
+Path signatures do not recover such latent states from CTR alone, but they can encode multiscale temporal structure in observed performance traces that may act as useful proxies for dynamics associated with fatigue.
 2.4 Change Point Detection: Methods and Limitations
@@
 (2020) providing comprehensive surveys of offline and online methods respectively.
-Classical approaches like CUSUM and MOSUM are computationally efficient but primarily designed for detecting changes in mean or variance, missing the complex dynamic changes that characterise creative fatigue.
+Common low-dimensional implementations of CUSUM and MOSUM are computationally efficient and often target changes in means or variances; they can miss richer trajectory dynamics unless the monitored statistic is designed to capture them.
 Killick et al.
-(2012) introduce the PELT algorithm for multiple change point detection with linear computational cost, representing a significant advance in scalability.
-However, their method still assumes piecewise stationarity, which may not hold for advertising data where performance can exhibit gradual degradation or volatile patterns.
+(2012) introduce the PELT pruning algorithm for multiple change point detection with linear computational cost under appropriate conditions, representing a significant advance in scalability.
+The modelling assumptions depend on the segment cost supplied to PELT; commonly used piecewise-constant or piecewise-stationary costs may be mismatched to advertising data where performance can exhibit gradual degradation or volatile patterns.
 Adams and MacKay (2007) propose Bayesian online change point detection that can handle more complex patterns, but at significant computational cost that limits scalability to large creative portfolios.
@@
 Path signature methods, rooted in rough path theory (Lyons, 1998), offer a fundamentally different approach to time-series analysis.
-Rather than focusing on statistical moments or model parameters, signatures provide a universal, non-parametric characterisation of a path’s geometry.
+Rather than focusing on statistical moments or model parameters, signatures provide a rich, non-parametric characterisation of a path’s geometry.
 Chevyrev and Kormilitzin (2016) provide an accessible introduction showing how signatures capture iterated integrals that encode rich information about path behaviour.
-The theoretical foundations established by Friz and Hairer (2014) demonstrate that signatures form a graded algebra that uniquely characterises paths up to tree-like equivalences, providing theoretical guarantees absent from many heuristic methods.
+The theoretical foundations established by Friz and Hairer (2014) demonstrate that signatures form a graded algebra, while uniqueness results for bounded-variation paths hold up to tree-like equivalence under appropriate hypotheses (Hambly and Lyons, 2010).
 Recent advances by Kidger et al.
 (2019) and Morrill et al.
-(2021) have shown that signature-based methods achieve state-of-the-art performance in various time-series tasks, from finance to healthcare.
-Fermanian (2021) establishes that signature embeddings preserve important geometric properties while providing dimensionality reduction, making them computationally tractable for real-time applications.
-This combination of theoretical rigour and practical efficiency makes signatures particularly suitable for marketing applications where both accuracy and scalability are critical.
+(2021) have shown that signature-based methods achieve strong performance in various time-series tasks, from finance to healthcare.
+Fermanian (2021) establishes useful properties of signature embeddings for learning from paths.
+In the present setting, the practical point is that a truncated signature gives a finite-dimensional fixed-depth embedding; it is not inherently a dimensionality reduction unless compared with a richer or longer path representation.
+This combination of theoretical rigour and fixed-depth tractability makes signatures attractive for marketing applications where auditability and scalability are important.
 2.6 Research Gap and Our Contribution
@@
 Existing approaches suffer from several limitations:
-First, traditional statistical methods (ARIMA, CUSUM) are designed for simple distributional changes and miss the complex dynamic patterns that characterise creative fatigue.
+First, simple traditional monitoring rules and common low-dimensional changepoint statistics are often tuned to specific distributional changes and may miss more complex dynamic patterns that characterise creative fatigue.
 Second, machine learning methods require extensive training data that may not exist for new creatives.
-Third, current methods lack the scalability required for modern programmatic advertising where thousands of creatives must be monitored simultaneously.
-Fourth, existing approaches provide binary detection without quantifying financial impact, limiting their utility for budget allocation decisions.
+Third, current methods can be difficult to audit at the creative level when many assets must be monitored simultaneously.
+Fourth, existing approaches often provide binary detection without a transparent prioritisation metric, limiting their utility for budget allocation discussions.
 Our contribution addresses these limitations by introducing path signatures to the marketing analytics domain.
-This geometric approach naturally captures complex temporal patterns without assuming specific functional forms, operates without training data, scales linearly with time-series length, and includes a novel framework for quantifying financial wastage.
-To our knowledge, this represents the first application of rough path theory to advertising analytics, opening new avenues for geometric methods in marketing science.
-The theoretical advantages of signatures—their universal approximation properties, invariance to time reparameterisation, and hierarchical structure—directly address the challenges of creative fatigue detection.
-By treating performance trajectories as geometric objects rather than statistical processes, we can detect subtle changes that traditional methods miss while maintaining the computational efficiency required for real-time application.
+This geometric approach captures temporal patterns without assuming a specific parametric functional form, can operate without labelled training data, scales linearly with time-series length for fixed signature depth, and includes a transparent framework for computing a benchmark-relative opportunity proxy.
+We position this as an early domain-specific application of rough path methods to advertising analytics and as a reproducible synthetic benchmark for future comparison.
+The theoretical advantages of signatures–their universal approximation properties, reparameterisation properties at the representation level, and hierarchical structure–make them attractive for screening candidate fatigue-related trajectory changes.
+These properties do not automatically transfer unchanged to the full detector, because the implemented pipeline embeds explicit time, uses fixed observed-day windows, and applies thresholding to finite samples.
+By treating performance trajectories as geometric objects rather than only as scalar sequences, the approach can detect changes in trajectory shape that simple level-based heuristics may miss, while retaining computational efficiency for fixed signature depth and window size.
 3 A Formal Framework for Fatigue Detection
@@
 where $\otimes$ denotes the tensor product.
-In our implementation, we compute these signatures using log-signatures (Lie increments) in the free Lie algebra, providing computational advantages while preserving the essential geometric information.
+In our implementation, signatures are computed using roughpy over increment streams; log-signature variants are natural alternatives but are not the default evaluated representation in this paper.
 The signature provides a non-parametric summary of the path.
-Chen’s theorem, a cornerstone of rough path theory, guarantees that the signature uniquely determines the path up to tree-like deformations, making it a powerful feature set.
-Before computing signatures, we apply a normalisation procedure to ensure comparability across different time windows.
+For bounded-variation paths with a fixed starting point, the full signature determines the path up to tree-like equivalence under appropriate hypotheses (Hambly and Lyons, 2010), making it a powerful feature set for path-shaped data.
+Before computing signatures, we apply a normalisation procedure to improve comparability across different time windows.
 Specifically, we scale time values to the interval $[0,1]$ and apply min-max scaling to metric values, also mapping them to $[0,1]$.
-This normalisation ensures that the signature captures the relative shape of the performance trajectory rather than its absolute scale, making the method robust to varying impression volumes and time spans.
-Within this framework, we can formally define a change point.
-We say that a change point $\tau$ occurs in the process $\{Y_{t}\}$ if the law of the process for $t>\tau$ is different from the law of the process for $t\leq\tau$.
-Our objective is to infer the existence of such a change point from the observed data.
-We frame this as a hypothesis testing problem.
-For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w_{d}$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
-$H_{0}:\text{Law}(Y_{t}|_{t\in W_{1}})=\text{Law}(Y_{t}|_{t\in W_{2}})$
-$H_{1}:\text{Law}(Y_{t}|_{t\in W_{1}})\neq\text{Law}(Y_{t}|_{t\in W_{2}})$
-Our test statistic for this hypothesis is the Euclidean distance between the normalised signatures of the paths corresponding to the two windows:
-$D=\|S(X_{W_{1}})-S(X_{W_{2}})\|_{2}$
-where $S(X_{W_{i}})$ denotes the truncated signature of the normalised path in window $W_{i}$.
-We chose the Euclidean distance for its computational efficiency and empirical effectiveness, though other metrics such as the signature kernel distance could be explored in future work.
-Under the null hypothesis, we expect this distance to be small, reflecting the random variation of a single, consistent process.
-Under the alternative hypothesis, we expect the distance to be large, reflecting a genuine change in the underlying dynamics of the path.
-A sufficiently large value of this test statistic leads us to reject the null hypothesis and declare the presence of a change point at the boundary between $W_{1}$ and $W_{2}$.
-4 Methodology: Signature-Based Change Point Detection
-Our proposed methodology for detecting creative fatigue is based on a systematic algorithmic process designed to identify statistically significant changes in the geometric properties of a performance time series.
-This section details the algorithm, the selection of its key parameters, and its computational complexity.
+When a metric window is constant, we map the metric coordinate to zero over that window; this makes the signature well-defined and treats flat windows as having no within-window metric movement.
+This normalisation induces a comparison of within-window relative shape, not the full law or absolute scale of the original process $Y_{t}$.
+All constant windows are indistinguishable after this transform, and pure level shifts between otherwise flat windows require a separate level feature, global scaling, or an exposure-aware representation.
+The transform also does not correct for heteroskedastic binomial noise caused by varying impression counts; impression-aware embeddings and exposure-weighted distances are therefore important extensions before production use.
+Classical changepoint terminology often defines a change point $\tau$ in a stochastic process $\{Y_{t}\}$ as a time where the law of the process after $\tau$ differs from the law before $\tau$.
+We use that notion only as background.
+The implemented detector targets a more local operational object: for a candidate boundary $\tau$, define two adjacent observed-day windows $W_{1}(\tau)=\{\tau-w_{d}+1,\ldots,\tau\}$ and $W_{2}(\tau)=\{\tau+1,\ldots,\tau+w_{d}\}$.
+A candidate change is flagged when these finite windows have unusually different path-shape summaries.
+Large local differences can arise from sampling variability as well as true process changes, so the detector should not be read as a test of a global pre/post law-change hypothesis.
+Let $\phi_{i}$ denote the window-specific transformation that rescales time to $[0,1]$ and min–max rescales the metric coordinate within window $W_{i}$, using a zero metric coordinate for flat windows.
+The transformed window path is $\widetilde{X}^{(i)}=\phi_{i}(X_{W_{i}})$.
+For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w_{d}$, we compute
+$D(\tau)=\|S(\widetilde{X}^{(1)})-S(\widetilde{X}^{(2)})\|_{2},$
+where $S(\widetilde{X}^{(i)})$ denotes the truncated signature of the transformed path in window $W_{i}$.
+This notation is important because signatures are not invariant to arbitrary coordinate scaling; the reported detector compares within-window relative shape rather than raw metric levels.
+We chose the Euclidean distance for its computational simplicity, though other metrics such as the signature kernel distance could be explored in future work.
+This score is a heuristic dissimilarity measure rather than a level-controlled statistical test: no null distribution or false-alarm level is claimed for the threshold used below.
+A sufficiently large value of $D(\tau)$ is treated as a candidate change point at the boundary between $W_{1}$ and $W_{2}$.
+4 Methodology: Level-Aware Signature-Based Change Point Detection
+Our proposed methodology for screening creative fatigue is based on a systematic algorithmic process designed to identify unusually large changes in the geometric properties of a performance time series while retaining the level deterioration central to fatigue.
+This section details the primary level-aware detector, the shape-only ablation, the selection of key parameters, and computational complexity.
+4.1 Primary level-aware detector
+The primary detector used for fatigue screening operates on the same noise-stabilised daily CTR scale used in the injected benchmark.
+Let $\widehat{p}_{t}$ be the Beta–Binomial posterior CTR estimate and let $\widetilde{CTR}_{t}$ be its trailing moving average.
+Using an initial calibration period $\mathcal{C}$, define
+$z_{t}=\frac{\widetilde{CTR}_{t}-\bar{c}_{\mathcal{C}}}{s_{\mathcal{C}}},$
+where $\bar{c}_{\mathcal{C}}$ and $s_{\mathcal{C}}$ are the mean and standard deviation of $\widetilde{CTR}_{t}$ in the calibration period, with a small positive floor on $s_{\mathcal{C}}$ for numerical stability.
+For each adjacent-window boundary $\tau$, construct two paths from $(u,z_{u})$ over the left and right windows after rescaling local time to $[0,1]$.
+The level-aware signature distance is
+$D_{\mathrm{level}}(\tau)=\|S(X^{L}_{\tau})-S(X^{R}_{\tau})\|_{2}.$
+In the past-only version, the threshold at boundary $\tau$ is computed from earlier distances only, after at least $h_{\min}$ earlier distances are available:
+$D_{\mathrm{level}}(\tau)>\mu_{<\tau}+k\sigma_{<\tau}.$
+A fatigue-specific alert additionally requires a directional deterioration gate,
+$\frac{\bar{\widetilde{CTR}}_{L(\tau)}-\bar{\widetilde{CTR}}_{R(\tau)}}{s_{\mathcal{C}}}>\gamma,$
+where $\gamma=2$ and $h_{\min}=25$ in the injected benchmark.
+This gate is not a signature term; it is an explicit domain restriction that distinguishes a fatigue screen from a generic trajectory-change detector.
+The exposure-aware variant augments the path to $(u,z_{u},\ell_{u})$, where $\ell_{u}$ is calibration-standardised $\log I_{u}$.
+Scoring boundary $\tau$ requires observing the full right-hand window, so the actionable online alert is available only at the end of that right window, not instantaneously at $\tau$.
+Algorithm 1 Past-Only Level-Aware Signature Fatigue Screen
+1:Input: Clicks $C_{t}$, impressions $I_{t}$, calibration period $\mathcal{C}$, window size $w_{d}$, depth $d$, threshold multiplier $k$, directional gate $\gamma$, minimum history $h_{\min}$.
+2:Compute posterior CTR estimates $\widehat{p}_{t}$ and smoothed series $\widetilde{CTR}_{t}$.
+3:Standardise $\widetilde{CTR}_{t}$ to $z_{t}$ using only $\mathcal{C}$.
+4:Initialise $Distances\leftarrow[]$ and $Alerts\leftarrow[]$.
+5:for each adjacent-window boundary $\tau$ after the calibration period do
+6:  Build left and right paths $X^{L}_{\tau},X^{R}_{\tau}$ from $(u,z_{u})$ with local time $u\in[0,1]$.
+7:  $D_{\tau}\leftarrow\|S(X^{L}_{\tau})-S(X^{R}_{\tau})\|_{2}$.
+8:  if DistancesDistances contains at least $h_{\min}$ past values then
+9:   $\theta_{\tau}\leftarrow\mu(D_{<\tau})+k\sigma(D_{<\tau})$.
+10:   $G_{\tau}\leftarrow(\bar{\widetilde{CTR}}_{L(\tau)}-\bar{\widetilde{CTR}}_{R(\tau)})/s_{\mathcal{C}}$.
+11:   if $D_{\tau}>\theta_{\tau}$ and $G_{\tau}>\gamma$ then
+12:     Append boundary $\tau$ and issue time $\tau+w_{d}-1$ to $Alerts$.
+13:   end if
+14:  end if
+15:  Append $D_{\tau}$ to $Distances$.
+16:end for
+17:return AlertsAlerts.
+4.2 Shape-only ablation
+The original shape-only detector uses independent min–max scaling in each window.
+It compares within-window trajectory shape and intentionally discards absolute level and amplitude differences.
+We retain it because it is useful for understanding what the signature representation alone contributes, but the injected benchmark shows that it is not sufficient for fatigue mechanisms dominated by level deterioration.
+Consequently, the level-aware detector above is the primary fatigue-screening method evaluated in the synthetic known-onset study.
 4.3 Algorithm and Pseudocode
-The core of our method is a four-step process that uses a sliding window to compare the geometric shape of a time series against itself.
+Algorithm 1 is the fatigue-specific online screen.
+The shape-only detector used in the panel experiment follows a related four-step process that uses a sliding window to compare the geometric shape of a time series against itself.
+It is included as an interpretable baseline and ablation rather than as the recommended fatigue-specific production rule.
 First, we employ a sliding window approach.
@@
 Second, for each window, we compute its corresponding path signature.
-The time-series data within the window is first normalised to ensure that the signature calculation is not sensitive to the absolute scale of the time or metric axes.
-Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$.
+The time-series data within the window is first normalised to reduce sensitivity to the absolute scale of the time or metric axes.
+Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$; flat metric windows are assigned a zero metric coordinate throughout.
 The normalised two-dimensional path is then passed to the library, which computes the path signature up to a pre-specified truncation depth $d$.
 This operation yields two signature vectors, $S(W_{1})$ and $S(W_{2})$, which serve as high-dimensional feature representations of the performance trajectory within each window.
-Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors.
-This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows.
-A large distance implies a significant change in the underlying dynamics of the time series.
-To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance.
-In our retrospective evaluation we compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series for a given creative.
-A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector.
-For online deployment, the same thresholding rule can be implemented using an initial calibration period or a rolling robust estimate of location and scale computed strictly from past distances.
+Because metric values are normalised within each window, this detector is intentionally invariant to absolute level and amplitude shifts; detecting those shifts requires a global scaling scheme or additional unnormalised features.
+Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D_{i}=\|S(W_{1}^{\text{norm}})-S(W_{2}^{\text{norm}})\|_{2}$, between the two signature vectors.
+This distance provides a scalar measure of dissimilarity in the transformed shape of the performance paths in the two adjacent windows.
+A large distance is treated as evidence of a candidate change in the underlying dynamics of the time series.
+To distinguish unusually large changes from background variation, we establish a threshold for this distance.
+In our retrospective panel evaluation we compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series for a given creative.
+This is an offline segmentation rule because early markers can depend on later distances.
+A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D_{i}$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector.
+This is an anomaly threshold, not a formal significance threshold.
+For online deployment, the same thresholding idea must be implemented using an initial calibration period or a rolling estimate of location and scale computed strictly from past distances.
 Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments.
 For each segment, we perform a final trend classification.
-We fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope.
+Segments with fewer than three observations are labelled “stable/insufficient data” because the usual slope significance test has no positive residual degrees of freedom, even though a two-point slope can be computed.
+For segments with at least three observations, we fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope.
 This final step provides an interpretable label for the performance behaviour within each period of stability, making the results directly actionable for a marketing analyst.
 The complete process is summarised in Algorithm 2.
-Algorithm 1 Signature-Based Change Point Detection
+Algorithm 2 Shape-Only Signature Change Point Detection
 1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w_{d}$, signature depth $d$, threshold multiplier $k$.
@@
 3:Initialise $Distances\leftarrow[]$.
-4:for $i\leftarrow 1$ to $T-2w_{d}+1$ do
-5:  $W_{1}\leftarrow Y[i:i+w_{d}]$
-6:  $W_{2}\leftarrow Y[i+w_{d}:i+2w_{d}]$
-7:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1], min-max scale metric to [0,1]
-8:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
-9:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
-10:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
-11:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
-12:  Append $(Y[i+w_{d}].\text{date},D_{i})$ to $Distances$.
-13:end for
+4:if $T<3$ then
+5:  return $[]$, “stable/insufficient data”
+6:end if
+7:if $T<2w_{d}$ then
+8:  return $[]$, $\text{ClassifyTrendViaRegression}(Y)$
+9:end if
+10:for $i\leftarrow 1$ to $T-2w_{d}+1$ do
+11:  $W_{1}\leftarrow Y[i:i+w_{d}-1]$
+12:  $W_{2}\leftarrow Y[i+w_{d}:i+2w_{d}-1]$
+13:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1]; min-max metric or use zero for flat windows
+14:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
+15:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
+16:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
+17:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
+18:  Append $(Y[i+w_{d}].\text{date},D_{i})$ to $Distances$.
+19:end for
 […]
@@
 […]
-20:for each SegmentSegment in SegmentsSegments do
-21:  $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
-22:  Append TrendTrend to $T_{s}$.
-23:end for
-24:return $CP,T_{s}$
+26:for each SegmentSegment in SegmentsSegments do
+27:  if $\text{Length}(Segment)<3$ then
+28:   $Trend\leftarrow\text{``stable/insufficient data''}$
+29:  else
+30:   $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
+31:  end if
+32:  Append TrendTrend to $T_{s}$.
+33:end for
+34:return $CP,T_{s}$
 4.4 Parameter Selection and Tuning
-The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w_{d}$ and the sensitivity threshold multiplier $k$.
-The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
+The performance of the signature-based change point detection algorithm is influenced by the window size $w_{d}$, signature depth $d$, threshold multiplier $k$, and, for the level-aware fatigue screen, the directional-drop gate $\gamma$.
+In the injected benchmark we use $d=3$, $w_{d}=14$, $k=1.5$ for the headline level-aware online row, and $\gamma=2$.
 The window size $w_{d}$ determines the granularity of the analysis.
@@
 A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events.
-The optimal choice of $w_{d}$ depends on the nature of the time series being analysed.
-For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
-The sensitivity threshold $k$ controls the trade-off between false positives and false negatives.
-It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change.
+The optimal choice of $w_{d}$ depends on the cadence and volatility of the time series being analysed.
+The experiments below use daily windows in the 7–14 day range as exploratory operating points, not as universal defaults.
+The sensitivity threshold $k$ controls the trade-off between false positives and false negatives through the rule
+$D_{i}>\mu_{D}+k\sigma_{D},$
+where $D_{i}$ is the adjacent-window signature distance and $\mu_{D},\sigma_{D}$ are computed from the relevant reference distribution of distances.
+In the retrospective panel experiment, the reference distribution is the full set of distances for that creative.
+In an online detector, it must be replaced by a past-only calibration distribution.
 A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives.
 A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes.
-A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data.
-The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
+In the experiments below, values between 1.5 and 2.5 are treated as exploratory operating points rather than universally recommended settings.
+The optimal choice should be tuned using validation data, explicit alert budgets, or historical data with known change points when available.
+For synthetic-only deployment research, we recommend a fixed calibration protocol: choose $(w_{d},d,k,\gamma)$ on a synthetic training suite before evaluating new scenarios.
+The objective should be stated before fitting, for example maximising
+$\text{mean captured proxy clicks per series}-c_{\mathrm{review}}\times\text{alert rate},$
+where alert rate is the fraction of series with at least one alert under the evaluated rule.
+Alternatively, one can target a no-event false-alert rate under nuisance regimes.
+In Section 7.2 we implement the simplest version of this protocol by selecting the threshold setting on the first 12 replicates per scenario and evaluating the frozen setting on the remaining 12 replicates.
+This does not replace external validation, but it prevents the headline threshold from being justified only by the same rows used for evaluation.
 4.5 Computational Complexity
@@
 Inside the loop, the most expensive operation is the signature computation.
-For fixed path dimension $m=2$ and small truncation depth $d\in\{3,4\}$, signature computation is effectively linear in the number of time points per window.
-Formally, the number of truncated signature terms scales as $\sum_{k=0}^{d}m^{k}$, so computational cost grows quickly with $d$ but remains practical for small $d$.
-Since $w_{d}$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$.
-This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
-5 Quantifying Wastage: The Financial Impact of Fatigue
-A key advantage of our methodology is its ability to translate a statistical observation into a tangible financial metric.
-While identifying a change point is analytically valuable, its true business utility is realised when we can quantify the cost of inaction.
-To this end, we introduce a framework for calculating ’wastage’, defined as the opportunity cost incurred by continuing to invest in a creative asset after its performance has begun to degrade.
-The calculation first requires the establishment of a fair performance benchmark.
-This benchmark represents the level of performance the creative was capable of achieving during its peak.
-We establish this by analysing the segments identified by the change point detection algorithm.
-The benchmark CTR, denoted $CTR_{bench}$, is calculated as the average CTR over the most effective period of stable or improving performance.
-This ensures the creative is judged against its own demonstrated potential rather than an arbitrary external standard.
-With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short.
-For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance.
-To reduce sensitivity to day-to-day sampling noise, we compute lost clicks using the same noise-robust CTR estimate used in the onset definition below.
-This is given by the formula
+Formally, the full truncated signature has $\sum_{k=0}^{d}m^{k}$ terms, equivalently $\sum_{k=1}^{d}m^{k}$ non-constant terms if the level-0 term is omitted by the implementation.
+For $N$ admissible window positions, the signature-computation term is approximately $\mathcal{O}(Nw_{d}\sum_{k=1}^{d}m^{k})$.
+Thus, the detector scales linearly in $T$ only when the window size, path dimension, and truncation depth are treated as fixed constants.
+This fixed-depth, fixed-window regime is the one used in our experiments and is suitable for portfolio screening; production deployment still requires end-to-end load testing.
+5 Quantifying Wastage: A Proxy for the Impact of Fatigue
+One practical reason to pair detection with valuation is that a statistical alert is useful only if it helps prioritise action.
+To this end, we introduce a benchmark-CPC proxy for ’wastage’, defined as the opportunity value of clicks not obtained when a creative performs below its own benchmark.
+This is not an identified causal estimate of financial loss; it is a consistent valuation rule for ranking and triage.
+The calculation first requires the establishment of a performance benchmark.
+For operational use, this benchmark must be computed using only data available up to the alert date, for example the best prior stable or improving segment identified before day $t$ or a detector-agnostic trailing recent-best period.
+The benchmark CTR, denoted $CTR_{bench}$, is defined on a comparable CTR scale using the same Beta-prior family as the daily estimator.
+One suitable definition is a posterior pooled benchmark over a benchmark period $B^{\star}$:
+$CTR_{bench}=\frac{\sum_{u\in B^{\star}}C_{u}+\alpha}{\sum_{u\in B^{\star}}I_{u}+\alpha+\beta}.$
+This impression-weighted construction aligns the benchmark with the subsequent lost-click calculation.
+Zero-impression days have no effect on the benchmark because they leave the pooled click and impression totals unchanged; the fixed prior contribution $\alpha+\beta$ is unaffected.
+The benchmark represents the level of performance the creative had recently demonstrated, not an externally imposed standard.
+With the benchmark established, we can calculate the opportunity proxy for any subsequent period where performance falls short.
+For each day $t$ after the benchmark period, we calculate a smoothed ’lost-click’ proxy.
+This is not the literal realised-click shortfall $CTR_{bench}I_{t}-C_{t}$; it is a noise-stabilised proxy used for ranking and comparison.
+To reduce sensitivity to day-to-day sampling noise, we compute it using the same noise-robust CTR estimate used in the onset definition below.
+Specifically, for daily clicks $C_{t}$ and impressions $I_{t}$ we use the posterior mean
+$\widehat{p}_{t}=\frac{C_{t}+\alpha}{I_{t}+\alpha+\beta},$
+where $\alpha=m_{0}\bar{p}$, $\beta=m_{0}(1-\bar{p})$, $m_{0}=1000$, and $\bar{p}$ is the pooled CTR across the synthetic panel.
+In an operational online workflow, $\bar{p}$ must be fixed from historical training data available before the scored period; in this paper it is an offline synthetic-panel construction.
+Smoothed lost clicks are then
 $\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-\widehat{p}_{t})\times\text{Impressions}_{t}$
-where $\widehat{p}_{t}$ is the Beta–Binomial shrinkage estimate of CTR and $\text{Impressions}_{t}$ is the number of impressions on day $t$.
-To translate these lost clicks into a monetary value, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period.
-The daily wastage is then
+where $\text{Impressions}_{t}$ is the number of impressions on day $t$.
+To translate these lost clicks into a monetary proxy, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period.
+The daily proxy is then
 $\text{Wastage}_{t}=\text{Lost Clicks}_{t}\times CPC_{bench}$
-The total wastage for a creative is the sum of this daily wastage over the entire period of underperformance.
-This framework provides a clear and interpretable financial metric that directly links the statistical detection of fatigue to a business-relevant outcome.
-It allows marketing managers to not only identify underperforming assets but also to prioritise them based on the financial impact of their decline, leading to more efficient and data-driven budget allocation.
+The total wastage proxy for a creative is the sum of this daily value over dates after the benchmark period for which $\widehat{p}_{t}<CTR_{bench}$.
+Alternative valuation rules could assign a different monetary value to lost clicks; the purpose here is to use a transparent and reproducible rule for comparing creatives and alert timings.
 6 Experimental Validation
 6.1 Synthetic Data Framework
-To address the confidentiality constraints inherent in digital advertising data while enabling methodological exposition and evaluation, we use a synthetic panel dataset designed to mimic common characteristics of performance marketing time series: (i) high-frequency daily measurement, (ii) substantial day-to-day volatility in CTR, and (iii) heterogeneous trajectories across creatives.
+To address the confidentiality constraints inherent in digital advertising data while enabling methodological exposition and evaluation, we use synthetic datasets only.
+This is a deliberate evidence boundary: the paper does not report client campaign histories, and it therefore does not claim external validation on production portfolios.
+The synthetic panel dataset is designed to mimic common characteristics of performance marketing time series: (i) high-frequency daily measurement, (ii) substantial day-to-day volatility in CTR, and (iii) heterogeneous trajectories across creatives.
 6.1.1 Dataset structure and provenance
 The synthetic dataset is included with the repository (data/synthetic/synthetic_data.csv).
-It contains 3,077 observations from 2025–01–01 to 2025–06–30 across 17 creatives spanning multiple campaigns.
-Each record contains impressions and clicks (allowing CTR to be treated as a noisy binomial rate), alongside auxiliary spend fields.
-Most creatives have near-daily coverage over the study window, but a subset are intentionally irregularly observed to reflect real-world missingness (e.g., pacing, pauses, or delivery gaps).
-The CSV also includes a pre-computed CTR column; throughout this paper we define $CTR_{t}:=C_{t}/I_{t}$ and use clicks and impressions for all modelling and evaluation.
+It contains 3,077 observations from 2025–01–01 to 2025–06–30 across 17 creatives spanning multiple campaigns, corresponding to a complete daily panel over this date range.
+Each record contains impressions and clicks, so for creative $i$ on observed day $t$ clicks can be modeled as $C_{it}\mid I_{it}\sim\mathrm{Binomial}(I_{it},p_{it})$, and the observed CTR $C_{it}/I_{it}$ is a noisy binomial proportion with variance decreasing in $I_{it}$.
+The CSV also includes a pre-computed CTR column; throughout this paper we define $CTR_{it}:=C_{it}/I_{it}$ for $I_{it}>0$ and use clicks and impressions for all modelling and evaluation.
 6.1.2 Alignment with empirical magnitudes
 The panel is calibrated to resemble plausible ranges reported in industry and marketing analytics contexts: impression volumes are typically in the thousands to tens of thousands per day, and CTR values concentrate in the low single-percent range (approximately 0.5%–3.5%), consistent with broad digital advertising benchmarks (IAB, 2023).
-The dataset is not intended to reproduce any specific platform or sector; rather, it provides a transparent setting for comparing detectors under a clearly specified operational ground truth.
+The dataset is not intended to reproduce any specific platform, sector, bidding system, targeting policy, or auction environment; rather, it provides a transparent setting for comparing detectors under a clearly specified rule-based evaluation target.
+The synthetic-only design should be read as proof-of-concept evidence rather than as a substitute for deployment validation.
 Importantly, the dataset does not include injected changepoints or labelled fatigue onsets.
 We therefore define fatigue onset operationally in Section 6.4 and evaluate methods against that rule-based definition.
+This design supports a proof-of-concept comparison against an explicit operational rule, not a claim of recovery against independently known fatigue events.
 6.2 Comparative Method Analysis
-We benchmark our signature-based approach against established methods from both marketing practice and statistical literature. Table 6.2 provides a comprehensive comparison:
+We benchmark our signature-based approach against established methods from both marketing practice and statistical literature.
+Table 6.2 provides an informal qualitative summary rather than a result established by the experiments in this paper:
 [hb]
-Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No Tunable High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(n\,w_{reg})$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
+Qualitative context for creative fatigue detection methods Method Cost Train Delay Robust.
+Interpret.
+Scale Path Signature $O(n)$ No Tunable High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(n\,w_{reg})$ No 2-4 days Low High Good PELT1 $O(n)$ No Offline segmentation Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
   • 1(Killick et al., 2012)
     2(Adams and MacKay, 2007)
-The signature-based approach offers several key advantages.
-Unlike machine learning methods, it requires no training and operates immediately on new creatives.
-Its responsiveness is tunable through the window size and threshold parameter, enabling a practical trade-off between early warning and alert burden.
-Its linear complexity enables scaling to portfolios containing thousands of creatives.
-Furthermore, the geometric robustness of the approach allows it to capture complex patterns beyond simple mean or variance shifts, making it particularly suited to the dynamic nature of advertising performance data.
+Table 6.2 is intended as qualitative context, not as an empirical ranking of all listed methods.
+The rule-based panel experiment evaluates the signature detector against two practical heuristic baselines because those baselines are common in operational monitoring.
+The injected-changepoint benchmark additionally includes MOSUM-style adjacent-window mean drops, a GLR-style adjacent-window mean-drop statistic, CUSUM, a hand-crafted window-feature detector, a Gaussian-kernel window-distance detector, and an offline mean-shift reference.
+A stronger validation would extend this comparison to PELT, Bayesian online changepoint detection, and additional heteroskedastic or exposure-weighted window methods under matched alert-burden constraints.
+The signature-based approach offers several practical advantages.
+Unlike supervised machine learning methods, it requires no labelled training set; however, in the main panel experiment its decision threshold is calibrated retrospectively rather than deployed prospectively on new creatives.
+Its responsiveness is tunable through the window size and threshold parameter, enabling a practical trade-off between earlier alerts and alert burden.
+Its fixed-depth complexity is suitable for portfolio screening.
+These advantages should be interpreted alongside the current experiment’s limitations: the evaluation target is rule-based, the main threshold is retrospective, and the present baseline set is not exhaustive.
 6.3 Experimental Design
@@
   2. 2.
-    Operational Ground Truth: Define fatigue onset $\tau$ per creative via the noise-robust, benchmark-relative rule in Section 6.4.
+    Rule-based evaluation target: Define fatigue onset $\tau$ per creative via the noise-robust, benchmark-relative rule in Section 6.4.
   3. 3.
-    Method Application: Apply the signature-based detector and two baseline heuristics (moving-average crossover; rolling regression) to each creative.
+    Method Application: In the panel experiment, apply the signature-based detector and two baseline heuristics (moving-average crossover; rolling regression) to each creative.
+In the injected benchmark, use the comparator set listed below.
   4. 4.
@@
   5. 5.
-    Sensitivity Analysis: Vary detector parameters to characterise the trade-off between lead time and alert burden.
+    Injected Benchmark: Run a stochastic simulation with known fatigue onset times and no-event controls, comparing retrospective and past-only signature variants against MOSUM-style mean drops, a GLR-style mean-drop statistic, CUSUM, an offline mean-shift baseline, an exposure-aware signature variant, a hand-crafted window-feature baseline, and a Gaussian-kernel adjacent-window distance baseline.
+  6. 6.
+    Ablation and stress tests: Vary signature depth and report false-alert rates on no-event nuisance scenarios, including impression shifts, seasonality changes, and delivery gaps without latent CTR fatigue.
+  7. 7.
+    Operational checks: Report a benchmark-relative lost-clicks proxy under each alerting rule and a runtime smoke benchmark for portfolio screening.
+  8. 8.
+    Sensitivity Analysis: Vary detector parameters to characterise the trade-off between retrospective timing and alert burden.
 For the baseline methods, we use $CTR_{t}:=C_{t}/I_{t}$ and define windows over observed days (i.e., the last $k$ available observations for a creative) to accommodate missingness.
 Moving-average crossover uses trailing simple moving averages $MA_{7}(t)$ and $MA_{21}(t)$ computed over the last 7 and 21 observed days, respectively.
+The 21-day moving average and rolling-regression baselines become eligible only once at least 21 observed days are available.
 An alert triggers on observed day $t$ if $MA_{7}(t)<MA_{21}(t)$ and this inequality holds for three consecutive observed days.
-Rolling regression fits an OLS model of $CTR$ on time over the last $w_{reg}=21$ observed days and triggers when the estimated slope is negative and statistically significant at $\alpha=0.05$ for three consecutive observed days.
-The experimental framework ensures fair comparison by: - Using identical data for all methods - Using consistent pre-processing and smoothing assumptions in the operational ground truth - Defining a common alert date $\hat{\tau}$ for each method as specified in Section 7 - Reporting sensitivity analysis over detector parameters to characterise the early-warning versus alert-burden trade-off
-6.4 Operational Definition of Fatigue Onset (Ground Truth)
+Rolling regression fits an unweighted OLS model of $CTR$ on time over the last $w_{reg}=21$ observed days and triggers when the estimated slope is negative and statistically significant at $\alpha=0.05$ for three consecutive observed days.
+This slope test is a simple heuristic baseline rather than a variance-calibrated test under the binomial noise model; exposure-weighted or heteroskedasticity-robust rolling regressions would be stronger comparators.
+The experimental framework supports within-experiment comparison by:
+  • using identical data for all methods within each experiment;
+  • using consistent pre-processing and smoothing assumptions in the rule-based evaluation target;
+  • defining a common timing marker $\hat{\tau}$ for each method as specified in Section 7; and
+  • reporting sensitivity analysis over detector parameters to characterise the early-warning versus alert-burden trade-off.
+6.4 Operational Definition of Fatigue Onset
 Our empirical evaluation requires a reproducible definition of “fatigue onset” that is robust to the substantial day-to-day volatility inherent in digital advertising metrics.
-Because the dataset does not contain injected changepoints or labelled onset dates, we adopt an operational ground truth based on a sustained, economically meaningful deterioration in click-through rate (CTR) relative to a recent performance baseline.
+Because the dataset does not contain injected changepoints or independently labelled onset dates, we adopt a rule-based evaluation target based on a sustained, economically meaningful deterioration in click-through rate (CTR) relative to a recent performance baseline.
+This target should not be read as external ground truth; it is an operational proxy used to make the comparison reproducible.
 Noise-robust CTR estimation.
-Let $I_{t}$ and $C_{t}$ denote daily impressions and clicks for a given creative, and define the observed CTR as $CTR_{t}=C_{t}/I_{t}$.
+Let $I_{t}$ and $C_{t}$ denote daily impressions and clicks for a given creative, and define the observed CTR as $CTR_{t}=C_{t}/I_{t}$ when $I_{t}>0$; zero-impression days are excluded from raw CTR calculations.
 In finite samples, $CTR_{t}$ exhibits non-trivial measurement noise (particularly when $I_{t}$ fluctuates).
@@
 We parameterise the prior by a notional prior sample size $m_{0}$ and a global mean CTR $\bar{p}$, setting $\alpha=m_{0}\bar{p}$ and $\beta=m_{0}(1-\bar{p})$.
-In our experiments, we use $m_{0}=1000$ impressions and estimate $\bar{p}$ as the pooled mean CTR across all creatives.
-We then define the smoothed CTR series as a trailing moving average of length $w_{s}$:
+In our experiments, we use $m_{0}=1000$ impressions and estimate $\bar{p}$ as the pooled mean CTR across all creatives, assuming $0<\bar{p}<1$.
+This pooled value is an offline evaluation device and means that one creative’s smoothed series depends weakly on the panel-level prior.
+An online deployment would need to fix $\bar{p}$ from a historical training baseline before scoring future days, or use a leave-one-creative-out or externally specified prior.
+We then define the smoothed CTR series as a trailing moving average of length $w_{s}$.
+For $t\geq w_{s}-1$,
 $\widetilde{CTR}_{t}=\frac{1}{w_{s}}\sum_{j=0}^{w_{s}-1}\widehat{p}_{t-j},$
-with $w_{s}=7$ to mitigate weekly noise while preserving responsiveness.
+and for initial times $t<w_{s}-1$ we use the available prefix,
+$\widetilde{CTR}_{t}=\frac{1}{t+1}\sum_{j=0}^{t}\widehat{p}_{t-j}.$
+We set $w_{s}=7$ to mitigate weekly noise while preserving responsiveness.
 Recent-best baseline and onset criterion.
-Fatigue is defined relative to a “recent best” baseline. For each day $t$, define
+Fatigue is defined relative to a “recent best” baseline. Assuming $1\leq w_{s}\leq B$, for each eligible day $t$ define
 $Base_{t}=\max_{u\in[t-B,\;t-w_{s}]}\widetilde{CTR}_{u},$
@@
 $Drop_{t}=1-\frac{\widetilde{CTR}_{t}}{Base_{t}}.$
-We declare that a creative has entered fatigue when the degradation exceeds a material threshold $\delta$ for a sustained period of $p$ consecutive days.
+The drop ratio is evaluated only when the baseline window is nonempty and $Base_{t}>0$.
+We declare that a creative has entered fatigue when the degradation exceeds a material threshold $\delta$ for a sustained period of $p$ consecutive observed days.
 Formally, the fatigue onset date $\tau$ is the earliest $t$ such that
-$Drop_{t^{\prime}}\geq\delta\quad\text{for all}\quad t^{\prime}\in\{t,t+1,\dots,t+p-1\}.$
-In the main specification, we set $\delta=0.15$ (a 15% sustained decline relative to recent-best performance) and $p=5$ days.
+$Drop_{s_{j+r}}\geq\delta\quad\text{for all}\quad r=0,\dots,p-1,$
+where $s_{j}$ denotes the ordered observed days and $\tau=s_{j}$.
+In the main specification, we set $\delta=0.15$ (a 15% sustained decline relative to recent-best performance) and $p=5$ observed days.
+The rule is based on observed-record persistence rather than calendar-day contiguity.
+Because the synthetic panel is complete, this distinction has no effect in the reported experiment.
+For sparse synthetic extensions, the rule should add a maximum-span constraint such as $s_{j+p-1}-s_{j}\leq G$ calendar days before declaring a sustained run.
 Creatives that never satisfy this criterion are treated as having no fatigue onset within the observed horizon.
 All quantities are evaluated only when the required history is available.
-For creatives with missing days, lookback windows are defined over the last $k$ observed days rather than calendar days.
-This onset definition is retrospective because it requires observing $p$ consecutive days; in an online setting, the earliest actionable alert corresponding to this definition would occur at day t+p1t+p-1.
+For creatives with missing days, let $s_{1}<\cdots<s_{n}$ be the observed days; then for $t=s_{j}$, compute $Base_{t}=\max\{\widetilde{CTR}_{s_{j-r}}:r=w_{s},\dots,B\}$ whenever those observed-day indices exist.
+This onset definition is retrospective because it requires observing $p$ consecutive qualifying observations; in an online setting, the earliest actionable alert corresponding to this definition would occur on the $p$-th qualifying observed day.
 Robustness checks.
-To assess sensitivity to the operational definition, we report robustness results over $\delta\in\{0.10,0.15,0.20\}$ and $p\in\{3,5,7\}$, holding the smoothing parameters fixed.
-This ensures that comparative conclusions are not artefacts of a single threshold choice.
+To assess sensitivity to the operational definition, the intended robustness check varies $\delta\in\{0.10,0.15,0.20\}$ and $p\in\{3,5,7\}$, holding the smoothing parameters fixed.
+The present draft reports the main operating point and detector-parameter sensitivity below; a complete submission should add the full $(\delta,p)$ grid for fatigue decisions, alert burden, and benchmark-relative loss estimates.
 6.5 Case Study Visualisations
-To demonstrate the practical application and interpretability of our methodology, we apply it to five synthetic datasets designed to mimic common real-world scenarios of creative performance decay.
-The resulting visualisations, shown in Figures 1 through 5, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns.
+To demonstrate the practical application and interpretability of our methodology, we apply it to seven synthetic case-study trajectories designed to mimic plausible scenarios of creative performance decay and missingness.
+The resulting visualisations, shown in Figures 1 through 7, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns.
 Each plot displays the daily CTR, the detected change points as vertical dashed lines, and the classified trend for each segment.
@@
 Figure 5: Multi-Stage Decline: A stepped decline with multiple distinct change points and periods of stability.
-Figure 6: Non-Continuous Data (Sharp Drop): Demonstrates the method’s robustness to significant gaps in the data.
-Figure 7: Non-Continuous Data (Gradual Decline): The signature method correctly identifies the change point despite irregular sampling.
-These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns.
-For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner.
-Figures 6 and 7 specifically illustrate that signature-based methods can be applied to irregularly sampled trajectories.
+Figure 6: Non-Continuous Data (Sharp Drop): Illustrates the detector on a trajectory with significant gaps in the data.
+Figure 7: Non-Continuous Data (Gradual Decline): The signature method detects a change point despite irregular sampling.
+These visualisations illustrate the algorithm on several fatigue-pattern examples.
+For a marketing analyst, these plots provide a narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner.
+Figures 6 and 7 provide qualitative examples under missingness; they are illustrative rather than a general robustness guarantee.
 Window-based baselines can also be applied under missingness, but require an explicit choice of whether windows are defined over calendar days or observed days.
@@
 The dataset contains 17 creatives over 2025–01–01 to 2025–06–30, of which 6 meet the operational fatigue onset criterion under the main specification ($\delta=0.15$, $p=5$, $w_{s}=7$, $B=21$, $m_{0}=1000$).
-For each method, we define an alert date $\hat{\tau}$ as the first alert date for that creative.
-For the signature-based approach, $\hat{\tau}$ is the start date of the first segment classified as “declining” (with a fallback to the first detected change point).
-For baselines, $\hat{\tau}$ is the first day on which their heuristic trigger persists for three consecutive observed days.
+For each forward-trigger baseline, we define an alert date $\hat{\tau}$ as the first calendar day on which that method’s alert condition is satisfied.
+If a method requires persistence for three observations, $\hat{\tau}$ is the first day of that qualifying run.
+For the signature-based approach in this panel experiment, we instead report a retrospective segment-start marker: the start date of the first segment classified as “declining” (with a fallback to the first detected change point).
+The signature threshold uses the full-series distribution of distances, so this marker can be earlier than the day on which an online detector could have generated an actionable alert.
+Online deployment requires a past-only calibration rule and should be evaluated separately.
 […]
-Table 1: Lead time and alert burden under an operational ground truth definition
-Method $N_{fat}$ $N_{hit}$ $N_{miss}$ $N_{burden}$ Median lead
+Table 1: Retrospective timing and alert burden under the rule-based onset definition
+Method $N_{fat}$ $N_{hit}$ $N_{miss}$ $N_{burden}$ Median signed timing $L$
 Path Signature ($w_{d}=14$, $k=2.5$, depth=4) 6 6 0 11 28
@@
 Table 1 highlights a practical trade-off.
-Both the signature-based detector and the moving-average crossover provide early warning for all fatigue creatives under the operational definition, but they also generate alerts for all non-fatigue creatives in this dataset.
-By contrast, the rolling regression baseline is conservative (low burden) but misses most fatigue onsets and is, on average, late when it does trigger.
-Large positive lead times should be interpreted cautiously: extremely early alerts can reflect a mismatch between a heuristic trigger and the operational onset definition, rather than “better” detection.
+The signature-based row marks a declining segment before the rule-based onset for all fatigue creatives, and the moving-average crossover triggers before the rule-based onset for all fatigue creatives under this definition.
+Both also flag all non-fatigue creatives in this dataset.
+At this operating point, they therefore provide no creative-level filtering on the panel experiment: every creative is flagged at least once.
+By contrast, the rolling regression baseline has zero non-fatigue creative-level burden in this table, but misses most fatigue onsets and is, on average, late when it does trigger.
+Large positive signed timing values should be interpreted cautiously: extremely early markers can reflect a mismatch between a heuristic trigger and the operational onset definition, rather than “better” detection.
+For the signature-based detector and the moving-average crossover, the current results motivate burden-constrained evaluation and stronger baselines rather than a claim of creative-level discrimination.
+7.2 Injected changepoint benchmark
+To separate agreement with the operational rule from recovery of known event times, we also run a stochastic benchmark with a deterministic injected changepoint at a known onset time.
+The benchmark contains 264 simulated creative trajectories of length 180 days: 72 event trajectories (abrupt drop, gradual decline, and volatile decline) and 192 no-event trajectories.
+The no-event regimes include stable controls, impression-volume shifts, seasonality changes, delivery gaps, targeting-mix shifts, auction-pressure shocks, attribution-reporting volatility, and creative-rotation cycles.
+Event scenarios have a known onset at day 90.
+Daily clicks are generated from a Binomial model conditional on impressions, with impression volumes drawn from a noisy weekly seasonal process.
+CTR is smoothed using the same Beta–Binomial shrinkage logic used in the rule-based panel evaluation.
+Because the detector is applied to Beta–Binomial smoothed CTR rather than latent CTR, impression-volume shifts can change the observed benchmark series through the shrinkage map even when latent click probability is unchanged.
+Some no-event regimes deliberately change observed or latent CTR for non-fatigue reasons; these cases test whether a detector over-interprets operational confounding as creative wear-out.
+Table 2: Synthetic injected-benchmark diagnostics
+Group Series Days Median impr. Impr. IQR CTR p10–p90
+Known-onset event 72 180 7358 4102 1.35–2.55%
+No-event control 192 180 7195 4862 1.49–2.05%
+Table 2 summarises the synthetic benchmark scale.
+It is included to make the synthetic-only evidence auditable: the simulations cover daily histories with thousands of impressions, low-single-digit observed CTR, and controlled no-event nuisance regimes, including simplified versions of several advertising-specific confounds.
+They still do not reproduce proprietary platform mechanics, auction systems, or human creative-review decisions.
+The results below should therefore be interpreted as internal validity and stress-test evidence under stated data-generating processes.
+Table 3: Synthetic scenario diagnostics around the injected reference day
+Scenario Type Latent CTR pre/post Observed CTR pre/post Impr. pre/post
+Stable No event 1.80/1.80% 1.81/1.80% 7040/7076
+Impression shift No event 1.80/1.80% 1.80/1.80% 8466/3842
+Seasonality shift No event 1.80/1.80% 1.80/1.80% 7704/7704
+Delivery gaps No event 1.80/1.80% 1.81/1.79% 7349/7418
+Targeting mix No event 1.80/1.45% 1.80/1.45% 8411/8550
+Auction pressure No event 1.80/1.80% 1.80/1.81% 7591/5460
+Attribution volatility No event 1.80/1.80% 1.76/1.84% 8355/8448
+Creative rotation No event 1.81/1.79% 1.81/1.80% 7956/8383
+Abrupt drop Event 2.40/1.50% 2.39/1.50% 7390/7528
+Gradual decline Event 2.40/1.66% 2.39/1.65% 8201/8328
+Volatile decline Event 2.40/1.50% 2.40/1.50% 7561/7611
+Table 3 gives a compact audit of the scenario mechanisms by comparing median pre/post quantities around the injected reference day.
+The event rows show deliberate post-onset degradation in latent CTR.
+The no-event nuisance rows are not all stationary: targeting mix and attribution volatility can move observed or latent CTR for reasons other than creative fatigue, while auction pressure, impression shifts, and delivery gaps primarily affect exposure.
+These rows are intentionally difficult controls; a useful screen should surface their risk separately rather than treating every post-reference deterioration as wear-out.
+For this benchmark, an event hit is defined as the first actionable alert falling in $\tau,\tau+45]$, where $\tau$ is the injected onset.
+For adjacent-window online methods, the scored boundary is between the left and right windows, but the actionable alert date is the final day of the right window because those observations are required to compute the statistic.
+Alerts before $\tau$ or after $\tau+45$ are counted as invalid event alerts, and any alert on no-event controls is counted as a no-event false alert.
+We evaluate the original shape-only signature detector separately from level-aware variants.
+The shape-only variant uses independent per-window min–max scaling and no directional mean-drop gate.
+The level-aware variants use calibration-standardised CTR so that absolute deterioration is retained, and require the post-change window mean to be at least two calibration standard deviations below the pre-change window mean.
+This separation makes clear which evidence supports the original geometric representation and which evidence supports the fatigue-specific level-aware extension.
+Table 4: Injected changepoint benchmark with nuisance controls
+Method Event hit (72) Median delay Invalid alert (72) No-event false (192) Series
+Signature shape-only 26% 15 72% 99% 72/192
+Level-aware sig. retros. 29% 6 71% 25% 72/192
+Level-aware signature 94% 6 6% 31% 72/192
+Level+exposure signature 96% 6 4% 32% 72/192
+Window features 94% 6 4% 33% 72/192
+Kernel window 93% 11 7% 47% 72/192
+GLR mean-drop 90% 6 10% 49% 72/192
+MOSUM mean-drop 50% 6 50% 93% 72/192
+CUSUM drop 78% 5 22% 57% 72/192
+Offline full-series ref. 100% 4 0% 12% 72/192
+Table 4 is deliberately conservative: pre-onset alerts are not credited as hits, and adjacent-window online methods are scored at the day when the right window has become observable.
+Under this scoring rule, the shape-only signature detector has limited value for a fatigue DGP dominated by level deterioration, while the primary level-aware online signature variant recovers most known events with lower no-event burden than several simpler online comparators.
+The hand-crafted window-feature baseline remains a close competitor, and the Gaussian-kernel and GLR-style baselines also recover many events but with higher nuisance burden.
+This is important: the benchmark supports the level-aware adjacent-window workflow, but it does not show that signatures dominate all non-signature window summaries.
+The MOSUM and GLR rows are intentionally strong online level-change comparators; their results help distinguish the value of the signature representation from the value of simply detecting adjacent-window mean deterioration.
+The offline full-series reference uses future observations and is not an operational online competitor; its row should be read as an oracle-style upper reference for this simplified one-change design.
+It performs best on this benchmark, which is expected because the data-generating process matches its assumptions more closely than it matches the broader dynamic-pattern objective of path signatures.
+This result narrows the claim: signatures are a useful screening statistic for some heterogeneous trajectories, but the present benchmark does not establish superiority over all strong feature-based changepoint screens.
+Table 5: Operating points for online detectors under matched scoring
+Method / setting Event hit Median delay Invalid event alert No-event false alert
+Signature shape-only ($k=1.0$) 8% 4 92% 100%
+Signature shape-only ($k=1.5$) 26% 15 72% 99%
+Signature shape-only ($k=2.5$) 18% 22 28% 56%
+Level-aware signature ($k=1.0$) 94% 6 6% 35%
+Level-aware signature ($k=1.5$) 94% 6 6% 31%
+Level-aware signature ($k=2.5$) 94% 6 4% 26%
+Level+exposure signature ($k=1.0$) 96% 6 4% 36%
+Level+exposure signature ($k=1.5$) 96% 6 4% 32%
+Level+exposure signature ($k=2.5$) 99% 6 1% 24%
+Window features ($k=1.0$) 92% 6 7% 39%
+Window features ($k=1.5$) 94% 6 4% 33%
+Window features ($k=2.5$) 83% 5 8% 24%
+Kernel window ($k=1.0$) 90% 10 10% 49%
+Kernel window ($k=1.5$) 93% 11 7% 47%
+Kernel window ($k=2.5$) 53% 13 3% 32%
+GLR mean-drop ($k=1.0$) 90% 6 10% 49%
+GLR mean-drop ($k=1.5$) 90% 6 10% 49%
+GLR mean-drop ($k=2.5$) 90% 6 10% 49%
+MOSUM mean-drop ($k=1.0$) 29% 5 71% 100%
+MOSUM mean-drop ($k=1.5$) 50% 6 50% 93%
+MOSUM mean-drop ($k=2.5$) 90% 6 10% 48%
+[…]
+[…]
+[…]
+Table 5 reports selected operating points under the same post-onset hit definition.
+The table makes the alert-budget trade-off explicit.
+For the level-aware variants, increasing the signature threshold lowers no-event false alerts but does not materially improve event recovery in this simulation.
+The shape-only rows are not monotone in event-hit rate because changing the threshold can move alerts from invalid pre-onset times into the post-onset scoring window; this reinforces that shape-only signatures are not the recommended fatigue screen.
+The kernel-window rows show that a generic distribution-distance comparator can recover many events, but its operating points are either slower and higher burden or substantially lower recall.
+The GLR-style mean-drop rows behave like a strong level-change screen with high recovery but higher nuisance burden than the level-aware signature rows.
+These operating curves are more informative than a single headline setting and support treating the method as a candidate screen whose threshold must be tuned to an explicit alert budget.
+Table 6: False-alert rates by no-event nuisance scenario for selected online methods
+Nuisance regime Shape Level Exp. Feat. Kernel GLR CUSUM
+Attribution volatility 100% 88% 75% 83% 92% 92% 100%
+Auction pressure 100% 4% 8% 4% 38% 38% 42%
+Creative rotation 100% 0% 0% 0% 0% 0% 4%
+Delivery gaps 96% 21% 25% 29% 50% 62% 67%
+Impression shift 100% 29% 25% 29% 42% 50% 50%
+Seasonality shift 100% 8% 17% 12% 25% 25% 46%
+Stable 100% 4% 12% 12% 29% 29% 50%
+Targeting mix 100% 96% 96% 92% 100% 100% 100%
+Table 6 reports false-alert rates separately by no-event nuisance scenario.
+The most difficult controls differ by method, but the broad pattern is that shape-only signatures are highly sensitive to non-fatigue trajectory changes, while level-aware variants reduce but do not eliminate nuisance alerts.
+This supports the ReviewBot concern that signatures are not, by themselves, a causal fatigue model.
+The kernel-window and GLR comparators are also sensitive to several nuisance regimes, especially targeting mix and attribution volatility, which reinforces that generic window dissimilarity or level-drop evidence is not enough without domain constraints.
+The exposure-aware signature variant adds useful information in some volume-driven regimes, but it remains sensitive to targeting, auction, attribution, and rotation mechanisms that are deliberately outside the fatigue event definition.
+Table 7: Signature-depth ablation for the past-only detector
+Depth Event hit Median delay No-event false alert
+2 94% 6 33%
+3 94% 6 31%
+4 96% 6 32%
+Table 7 shows that increasing depth from 2 to 4 does not materially improve this benchmark.
+This is useful negative evidence: for the simulated fatigue mechanisms considered here, most of the recoverable signal is already available at low signature depth.
+More complex creative trajectories may require higher-depth or log-signature features, but the current evidence does not depend on depth 4.
+Table 8: Component ablation for the past-only level-aware detector
+Variant Event hit Median delay Invalid alert No-event false alert
+Shape-only signature 26% 15 72% 99%
+Directional gate only 90% 6 10% 49%
+Level-aware sig., no gate 33% 2 67% 96%
+Level-aware sig. + gate 94% 6 6% 31%
+Level+exposure sig. + gate 96% 6 4% 32%
+Table 8 separates the main ingredients of the level-aware detector.
+The directional gate alone is a simple mean-drop screen; the level-aware signature without the gate measures unusual adjacent-window path changes but does not enforce fatigue direction.
+The combined level-aware signature plus gate is the primary candidate rule, while the exposure-aware row adds standardised log-impressions as an additional path coordinate.
+In this benchmark, the gate supplies most of the event recovery, while adding the signature threshold reduces no-event burden relative to the gate-only rule.
+Exposure augmentation does not eliminate nuisance false alerts.
+The ablation therefore supports a restrained interpretation: signatures contribute a structured trajectory-change score, but the domain-specific deterioration gate carries important practical content.
+Table 9: Benchmark-relative impact proxy after detector alert
+Method Median missed clicks Post-onset loss flagged
+Signature shape-only 0 100%
+Level-aware sig. retros. 0 100%
+Level-aware signature 227 95%
+Level+exposure signature 231 95%
+Window features 231 95%
+Kernel window 334 92%
+GLR mean-drop 214 96%
+MOSUM mean-drop 0 100%
+CUSUM drop 136 97%
+Offline full-series ref. 195 96%
+Table 9 connects alert timing to the benchmark-relative lost-clicks proxy from Section 5.
+The table reports the median lost clicks that occur after the injected onset but before the first detector alert, along with the median share of post-onset loss covered by the alert timing.
+The summary includes all event trajectories: pre-onset alerts have zero post-onset missed loss, while no-alert trajectories have all post-onset loss counted as missed.
+A high captured-loss value can therefore be produced by over-early or over-frequent alerts, so the invalid-alert and no-event burden columns in Table 4 must be read alongside this proxy table.
+These are not causal financial estimates; they are a transparent way to compare prioritisation consequences under different alert rules.
+Table 10: Cost-sensitive operating-point selection on injected benchmark
+Method / selected setting Event hit No-event false alert Alert rate Captured clicks Net clicks
+Level-aware signature ($k=2.0$) 96% 28% 48% 1389 1318
+Level+exposure signature ($k=2.5$) 99% 24% 45% 1384 1317
+[…]
+Window features ($k=1.5$) 94% 33% 51% 1382 1306
+GLR mean-drop ($k=3.0$) 90% 48% 62% 1394 1301
+MOSUM mean-drop ($k=2.5$) 90% 48% 62% 1391 1297
+Signature shape-only ($k=1.0$) 8% 100% 100% 1431 1281
+Kernel window ($k=1.0$) 90% 49% 63% 1365 1271
+Table 10 adds a stylised operating-point exercise.
+For each online method family, it selects the threshold with the highest mean net proxy clicks per simulated series after charging a review cost of 150 clicks for every alert.
+The objective is
+[…]
+where alert rate is the fraction of simulated series producing at least one review.
+Net clicks are computed from unrounded alert rates and captured-click values; percentages in the table are rounded for display.
+The units are click-equivalent proxies rather than financial estimates.
+This table gives a concrete way to choose a threshold when analyst attention is costly: a method with high event recovery can still be unattractive if it creates many low-value reviews.
+Table 11: Train-selected thresholds evaluated on held-out synthetic replicates
+Method / frozen setting Train net Holdout hit Holdout false alert Alert rate Holdout net
+[…]
+GLR mean-drop ($k=3.0$) 1353 86% 49% 63% 1248
+Level+exposure signature ($k=3.0$) 1379 94% 25% 45% 1245
+MOSUM mean-drop ($k=2.5$) 1350 89% 52% 65% 1244
+Window features ($k=1.5$) 1369 89% 40% 55% 1242
+Level-aware signature ($k=3.0$) 1378 89% 26% 45% 1233
+Kernel window ($k=1.0$) 1330 86% 51% 64% 1212
+Table 11 turns the same idea into a train/held-out protocol.
+Thresholds are selected on the first 12 replicates of each synthetic scenario using the net-click objective above, then frozen and evaluated on the remaining 12 replicates.
+The level-aware signature setting remains competitive with the feature, kernel, and GLR baselines under this synthetic split and produces a lower no-event false-alert rate than the CUSUM, GLR, and kernel settings selected by the same objective.
+Because the split is over replicates from the same scenario families, this table should be read as a leakage-control check, not as evidence that the synthetic nuisance library is complete.
+Table 12: Fixed-capacity alert triage by alert-time opportunity proxy
+Method Capacity Reviewed Precision Post-onset loss captured
+Level-aware signature 10 10 100% 18%
+Level-aware signature 20 20 85% 31%
+Level+exposure signature 10 10 100% 18%
+Level+exposure signature 20 20 90% 33%
+Window features 10 10 100% 18%
+Window features 20 20 85% 31%
+Kernel window 10 10 100% 19%
+Kernel window 20 20 80% 27%
+GLR mean-drop 10 10 100% 18%
+GLR mean-drop 20 20 80% 29%
+CUSUM drop 10 10 70% 13%
+CUSUM drop 20 20 65% 23%
+Table 12 reframes the same alerts as a finite analyst-capacity problem.
+Within each detector’s alerts, candidate reviews are ranked by the alert-time lost-click proxy computed from the calibration benchmark and the current posterior CTR.
+The table reports the top 10 and top 20 reviews for selected online methods.
+Precision is the share of reviewed alerts that fall inside the post-onset hit window, while loss captured is the share of total injected post-onset loss associated with those reviewed alerts.
+This is a triage diagnostic rather than a causal value estimate, but it makes the practical capacity trade-off more explicit than detector hit rates alone.
+Table 13: Sensitivity of opportunity ranking to benchmark definition
+Benchmark rule Median lost clicks Rank corr. Top-10 overlap
+Calibration mean 5097 1.00 100%
+Calibration 75th pct. 5687 1.00 90%
+Pre-onset recent best 6306 0.98 90%
+Table 13 checks whether the opportunity ranking is stable to reasonable benchmark definitions.
+In this injected benchmark, the rank correlation with the calibration-mean benchmark remains high and the top-10 overlap is at least 90%, although the median lost-click magnitude rises when the benchmark uses a higher pre-onset reference.
+This supports the proxy as a ranking aid in the synthetic setting, while showing that absolute loss magnitudes remain benchmark-dependent.
+Table 14: Worked toy signature distances for simple CTR paths
+Contrast Same endpoints Depth 1 Depth 2 Depth 3
+monotone vs zig-zag Yes 0.00000 0.00000 0.00055
+monotone vs recover/fall Yes 0.00000 0.00707 0.00791
+monotone vs flat No 0.01000 0.01225 0.01258
+Table 14 gives a small representation sanity check.
+The monotone and zig-zag paths have the same start and end CTR, so their depth-1 displacement distance is zero.
+Higher-depth terms can distinguish some equal-endpoint path shapes, as seen in the depth-3 distance for the zig-zag contrast and the depth-2/3 distances for the recover-then-fall contrast.
+This toy example illustrates the geometric contribution of signatures, but it is not by itself evidence of superior detector performance under noisy advertising data.
+Table 15: Worked online level-aware detector trace
+Case Boundary Available Raw CTR L/R $\widetilde{CTR}$ L/R $D$ $\theta$ $G$ / decision
+Abrupt drop 84 97 2.31/1.88% 2.30/2.14% 161.98 34.77 2.02 / alert
+Targeting mix 85 98 1.84/1.59% 1.81/1.69% 82.04 51.36 2.16 / false alert
+Stable 100 1.79/1.78% 1.80/1.76% 7.71 7.36 0.88 / no alert
+Table 15 walks through the primary online detector on three benchmark cases.
+The boundary column is the adjacent-window split being scored; the available column is the first day on which the alert can be issued after observing the right window.
+The abrupt-drop row shows the intended mechanism: raw and smoothed CTR deteriorate, the signature distance exceeds its past-only threshold, and the directional gate is satisfied.
+The targeting-mix row is deliberately included as a failure mode: the same conditions can be met by a non-fatigue regime shift, so the method should be treated as a screen for analyst review rather than a causal fatigue classifier.
+The stable row shows a no-alert case where the displayed boundary is the closest eligible score to a joint threshold crossing.
+Table 16: Runtime smoke benchmark for past-only signature detection
+Series Depth Seconds Projected sec. / 1000 series
+25 3 0.54 21.5
+25 4 0.64 25.5
+100 3 2.14 21.4
+100 4 2.51 25.1
+Finally, Table 16 provides a small runtime smoke benchmark for past-only signature detection on replicated synthetic series using full-path roughpy signatures.
+The results are intended to validate order-of-growth plausibility, not to replace an end-to-end system benchmark.
+They suggest that fixed-depth portfolio screening is computationally feasible at the scale of thousands of daily creative trajectories on the test machine, while a deployed system would still need timing that includes data loading, report generation, storage, and alert delivery.
+7.3 Synthetic-only evidence boundary
+No client campaign histories are used in this paper.
+This avoids confidentiality risk and keeps the results fully reproducible, but it also limits the claims.
+The experiments can show whether the proposed detectors behave sensibly under transparent synthetic data-generating processes with known onsets and nuisance controls.
+They cannot establish that the detector will identify fatigue in a particular production portfolio, that the synthetic nuisance set spans all operational confounders, or that the benchmark-relative opportunity proxy will match analyst-reviewed business value.
+For this reason, the contribution is framed as a synthetic evaluation of a screening framework rather than as a production-calibrated fatigue detector.
+Under a synthetic-only evidence standard, the main quality controls are specification, coverage, and reproducibility.
+The data-generating mechanisms, onset scoring rule, no-event scenarios, detector variants, and generated tables are all included in the repository under presentations/paper/experiments.
+Future synthetic extensions should deepen the nuisance library further to include platform learning phases, intervention feedback loops, multi-stage creative deterioration, and cross-campaign dependence while preserving the same aggregate reporting discipline.
 7.4 Sensitivity analysis for the signature-based detector
-We next vary the signature detector’s window size $w_{d}$ and threshold multiplier $k$ to characterise the lead-time versus alert-burden trade-off.
-Table 2 reports the number of fatigue creatives alerted (out of 6), burden (out of 11), and the median lead time over fatigue creatives.
+We next vary the signature detector’s window size $w_{d}$ and threshold multiplier $k$ to characterise the alert-timing versus alert-burden trade-off.
+Table 17 reports the number of fatigue creatives alerted (out of 6), burden (out of 11), and the median signed detection offset relative to the fatigue reference date over alerted fatigue creatives only.
+The offset is $L=\tau-\hat{\tau}$, so positive values indicate a marker before the reference date and negative values indicate a marker after it.
 Table 17: Sensitivity of signature-based detection to $w_{d}$ and $k$ (depth=4)
-Window $w_{d}$ Threshold $k$ Alerts on fatigue Burden / Median lead (days)
+Window $w_{d}$ Threshold $k$ Alerts on fatigue Burden / Median signed $L$
 7 1.5 6/6 11 / 66.5
@@
 7.4.1 Financial Impact Quantification
-The true value of earlier detection becomes apparent when translated to operational units.
+The value of earlier review can be translated into operational units.
 The following calculation is a stylised illustration (not an empirical estimate from our synthetic panel).
-Given daily impressions $I$, benchmark cost-per-click $CPC_{bench}$, and an absolute CTR shortfall of $\Delta CTR$ relative to a benchmark period, the implied opportunity cost per day is
-$\text{Opportunity Cost per day}=I\,\Delta CTR\,CPC_{bench}.$
+Given daily impressions $I$, benchmark cost-per-click $CPC_{bench}$, and a nonnegative CTR shortfall $\Delta CTR\geq 0$ relative to a benchmark period, the implied spend-equivalent proxy per day is
+$\text{Spend-equivalent proxy per day}=I\,\Delta CTR\,CPC_{bench}.$
 This mapping is provided only to translate lead-time into operational units; it is not intended as a validated estimate of platform spending dynamics.
@@
 The signature-based detection algorithm is designed to be integrated into existing marketing analytics workflows.
-In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series, as discussed in Section 4.
-This makes it highly scalable and efficient.
-A typical analysis for a single creative with a year’s worth of daily data can be completed in a fraction of a second on standard hardware.
-The methodology can be deployed in either a batch or a real-time setting.
-In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an entire portfolio of active creatives, generating a summary report of assets that may be exhibiting fatigue.
-In a more advanced, real-time implementation, the algorithm could be integrated with a streaming data pipeline, allowing for the continuous monitoring of creative performance and the generation of automated alerts when a potential change point is detected.
-This would enable the fastest possible response to creative fatigue.
+In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series for fixed path dimension and signature depth, as discussed in Section 4.
+The runtime smoke benchmark in Table 16 supports the feasibility of batch portfolio screening on synthetic trajectories, although it does not replace an end-to-end system benchmark.
+The methodology is most directly supported as a batch or periodic screening tool.
+In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an active creative portfolio, generating a summary report of assets that may be exhibiting fatigue.
+A deployable online protocol should be specified before operational use.
+A conservative version is: require at least 45 observed calibration days, compute signatures on the Beta–Binomial smoothed CTR series, estimate the distance threshold only from past distances, require a directional post-window mean drop for fatigue-specific alerts, suppress repeat alerts for a fixed cooldown period, and tune the threshold using an explicit review-cost or alert-budget objective such as Table 10.
+The pure shape-only signature detector can still be useful as a generic trajectory-change screen, but the injected benchmark indicates that fatigue monitoring should use the level-aware rule.
+Real-time implementation would still require exposure-aware noise handling, alert-budget controls, and portfolio-level runtime validation before deployment claims could be made.
 8.2 Managerial Implications
-The primary implication of this work for marketing managers is the ability to move from a reactive, heuristic-based approach to creative management to a proactive, data-driven one.
-By providing a reliable and early warning system for creative fatigue, our methodology enables several key managerial actions.
-First, it allows for more efficient budget allocation.
-Instead of continuing to spend on a creative that is no longer effective, managers can reallocate that budget to higher-performing assets or to the development of new creative.
-The wastage calculation provides a direct financial justification for these decisions.
+The primary implication of this work for marketing managers is a more structured screening input for periodic creative review, rather than a fully validated replacement for existing heuristic decision rules.
+By providing a screening signal for possible creative fatigue, our methodology enables several managerial actions.
+First, it can support more efficient budget allocation discussions.
+Instead of relying only on manual review or monthly averages, managers can inspect creatives whose trajectories have changed and decide whether they warrant further diagnosis before any budget reallocation or creative refresh.
+The wastage proxy provides a prioritisation signal, not a direct causal financial justification.
 Second, it provides a quantitative basis for the creative refresh cycle.
-The output of the analysis can be used to prioritise which creatives need to be replaced or updated most urgently, based on the magnitude of their performance decline and the associated financial wastage.
-Third, it creates a more rigorous feedback loop for the creative development process itself.
-By analysing which types of creatives fatigue more quickly, marketing teams can gain insights into what resonates most with their audience over the long term, informing the development of more durable and effective advertising campaigns in the future.
+The output of the analysis can be used to prioritise which creatives should be reviewed most urgently, based on the magnitude of their performance decline and the associated benchmark-relative opportunity proxy.
+Third, it can create a more rigorous feedback loop for the creative development process itself.
+By tracking which types of creatives are repeatedly flagged for review, marketing teams can generate hypotheses about durability and audience response that should be checked against campaign context before being treated as substantive audience-preference findings.
 9 Ethical Considerations and Reproducibility
 9.1 Ethical Considerations
-The methodology presented in this paper operates on aggregated, anonymised performance data.
-It does not require access to any personally identifiable information (PII) or individual user-level data, and as such, it aligns with modern data privacy principles such as those outlined in GDPR and other regulations.
+The methodology presented in this paper can operate on aggregated performance data and does not require individual user-level data.
+This is privacy-favourable compared with user-level tracking, but it does not by itself establish GDPR compliance; compliance depends on the broader data collection, consent, retention, and governance context in which the method is deployed.
 The analysis focuses on the performance of the creative asset itself, not on the behaviour of individual users.
@@
 10 Conclusion
-In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures.
-Our findings from a synthetic panel evaluation indicate that this signature-based approach can provide early warnings under an explicit operational onset definition, while highlighting a practical trade-off between early warning and alert burden.
-By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect changes in the underlying dynamics of creative performance that simpler models may miss.
-The subsequent quantification of this fatigue in terms of opportunity cost provides a bridge from statistical outputs to operational prioritisation.
-The practical implications of this methodology for marketing practitioners are significant.
-The automated and scalable nature of the algorithm allows for the continuous monitoring of a large portfolio of creative assets, something that is infeasible to do manually.
-This enables marketing and media teams to move from a reactive to a proactive stance on creative management.
-Rather than waiting for performance to decline substantially, they can be alerted to the onset of fatigue in near real-time.
-The wastage metric provides a clear, data-driven rationale for when to retire an underperforming creative and reallocate its budget, thereby improving the overall efficiency and return on investment of advertising campaigns.
+In this paper, we have introduced a methodology for screening advertising creative fatigue by applying the mathematical framework of path signatures.
+Our findings from a synthetic panel evaluation and stochastic injected-changepoint benchmark indicate that this signature-based approach can provide a structured trajectory-change signal, while highlighting a practical trade-off between event recovery and alert burden.
+In the injected benchmark, the original shape-only signature detector is not sufficient for fatigue dominated by absolute level deterioration; within that benchmark, the most defensible candidate workflow among the evaluated online signature variants is the level-aware signature screen with explicit directional gating and threshold tuning.
+The subsequent benchmark-relative opportunity proxy provides a bridge from statistical outputs to operational prioritisation.
+The practical implication of this methodology is a structured screening workflow for creative review.
+The detector can highlight candidate creatives whose performance trajectory has changed, and the benchmark-relative wastage proxy can help prioritise analyst attention.
+The present synthetic evidence does not show that the method should autonomously retire creatives or reallocate budgets; rather, it supports use as one input into a broader decision process.
+Future work within the same data-governance constraint should extend the baseline set further (for example PELT, Bayesian online changepoint detection, and exposure-weighted changepoint rules), evaluate broader synthetic nuisance scenarios, report full alert-budget curves, and improve synthetic calibration diagnostics.
 Appendix A A Deeper Dive into Rough Path Theory
 Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems.
-At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path.
-For a continuous path $X_{t}$ in $\mathbb{R}^{m}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{m}$.
-The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path.
-$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}$
+At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of an appropriately regular path.
+For a bounded-variation path $X_{t}$ in $\mathbb{R}^{m}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{m}$.
+More irregular continuous paths require an appropriate rough-path enhancement before these iterated integrals are well defined.
+The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path. For bounded-variation paths,
+$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}=X_{1}^{i}-X_{0}^{i}.$
 The second level, $S^{2}(X)$, is a matrix whose entries are iterated integrals.
 $S^{2}(X)_{ij}=\int_{0}^{1}\int_{0}^{t}dX_{s}^{i}dX_{t}^{j}$
-These second-order terms capture the area swept out by the path in the $ij$-plane. Higher-order terms capture more complex geometric properties.
-A key result from rough path theory, Chen’s theorem, states that the signature uniquely determines the path up to tree-like deformations.
-This means that if two paths have the same signature, they are essentially equivalent from a geometric perspective.
-This uniqueness property is what makes the signature such a powerful feature set for describing path-like data.
-In practice, we use a truncated signature, which provides a finite-dimensional approximation that is sufficient for most machine learning applications.
-Our library provides an efficient and numerically stable implementation for computing these truncated signatures.
+These second-order terms capture ordering information in the path increments.
+The antisymmetric component, $\frac{1}{2}(S^{2}_{ij}-S^{2}_{ji})$, corresponds to signed oriented area in the $(i,j)$ plane; the full second level should not be interpreted simply as planar area.
+Higher-order terms capture more complex geometric properties.
+Chen’s identity states that signatures are multiplicative under path concatenation, which is central to their algebraic usefulness.
+Separately, uniqueness results for bounded-variation paths with a fixed starting point show that the full signature determines the path up to tree-like equivalence under suitable conditions (Hambly and Lyons, 2010).
+In practice, we use a truncated signature, which provides a finite-dimensional approximation to this infinite object.
+Truncation generally loses injectivity: many distinct paths can share the same low-order signature terms.
+The empirical question is therefore whether the chosen truncation depth preserves the path features relevant to advertising-performance monitoring.
 Appendix B Algorithm Parameters
@@
 A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events.
-The optimal choice of $w$ depends on the nature of the time series being analysed.
-For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
-The sensitivity threshold $k$ controls the trade-off between false positives and false negatives.
-It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change.
+The optimal choice of $w$ depends on the cadence, volatility, and review objective of the time series being analysed.
+The experiments in this paper use daily windows in the 7–14 day range as exploratory operating points, not as universal defaults.
+The sensitivity threshold $k$ controls the trade-off between false positives and false negatives through $D_{i}>\mu_{D}+k\sigma_{D}$, where $\mu_{D}$ and $\sigma_{D}$ are computed from the chosen reference distribution of adjacent-window distances.
+In the retrospective panel experiment this distribution is the full creative-specific distance series; in an online detector it must be a past-only calibration distribution.
 A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives.
 A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes.
-A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data.
-The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
-Appendix C Mathematical Foundations and Theoretical Analysis
+Values between 1.5 and 2.5 are explored in this paper as operating points, not as universal defaults.
+The optimal choice can be fine-tuned using a validation set, alert budget, or historical data with known change points if available.
+Appendix C Mathematical Foundations and Theoretical Remarks
 This section provides informal theoretical support for why signature distances are sensitive to changes in the underlying trajectory.
-For readability, we present proof sketches rather than fully detailed proofs.
+The statements below are not used as formal guarantees for the empirical results.
+They are intended to clarify the intuition behind the representation and to state the assumptions that would be needed for stronger theory.
 We assume paths are continuous and of bounded variation (or, more generally, finite $p$-variation) so that truncated signatures are well-defined.
@@
 C.1 Signature distance sensitivity
-Theorem 1 (Signature sensitivity to path changes (proof sketch)).
-Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$.
-If there exists a time interval $t_{1},t_{2}]\subset[0,T]$ where the paths exhibit different geometric properties (trend, volatility, or curvature), then:
-$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|,$
-where $C>0$ depends on regularity assumptions and the truncation depth $d$.
-Proof.
-The signature captures iterated integrals of the path increments. For the first level:
+The signature captures iterated integrals of path increments. For the first level:
 $S^{1}(X)_{[t_{1},t_{2}]}-S^{1}(Y)_{[t_{1},t_{2}]}=\int_{t_{1}}^{t_{2}}(dX_{t}-dY_{t})=X(t_{2})-X(t_{1})-(Y(t_{2})-Y(t_{1}))$
-This directly captures the difference in total displacement. For the second level, which captures signed area:
+This directly captures differences in total displacement. For the second level:
 $S^{2}(X)_{[t_{1},t_{2}]}^{ij}-S^{2}(Y)_{[t_{1},t_{2}]}^{ij}=\int_{t_{1}}^{t_{2}}\int_{t_{1}}^{t}(dX_{s}^{i}dX_{t}^{j}-dY_{s}^{i}dY_{t}^{j})$
-This term is sensitive to differences in the path’s oscillatory behaviour and volatility. By the Chen-Strichartz formula, the signature norm satisfies:
-$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|\geq\|X(T)-X(0)-(Y(T)-Y(0))\|$
-For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound.
-A rigorous statement requires explicit norms and assumptions (e.g., bounded $p$-variation) and follows from continuity estimates for truncated signatures.
-∎
-C.2 Theoretical Analysis of Detection Power
-Theorem 2 (Detection power (heuristic bound)).
-Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
+For multidimensional embeddings, second-level cross terms encode ordering information through interactions between channels.
+In one-dimensional paths, the second level reduces to a function of endpoint displacement,
+$S^{2}(X)^{11}_{[t_{1},t_{2}]}=\frac{1}{2}\{X(t_{2})-X(t_{1})\}^{2},$
+and therefore does not by itself capture oscillation.
+If the truncated signature is equipped with a Euclidean direct-sum norm whose level-1 block has unit weight, the distance dominates the first-level displacement difference:
+$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|=\|X(T)-X(0)-(Y(T)-Y(0))\|.$
+Under other equivalent finite-dimensional norms, the same statement holds only up to a norm-dependent constant.
+Higher-order and cross-channel terms can distinguish some paths with similar endpoints but different shapes, but a general deterministic lower bound in terms of the path sup-norm requires additional assumptions and is not claimed here.
+C.2 Depth-1 relation to adjacent-window statistics
+It is useful to spell out what the level-aware construction reduces to at depth 1. Consider two adjacent windows represented as paths
+$X^{L}(u)=(u,z_{L}(u)),\qquad X^{R}(u)=(u,z_{R}(u)),\qquad u\in[0,1],$
+where $z$ is the calibration-standardised smoothed CTR. Their first-level signatures are endpoint increments:
+$S^{1}(X^{L})=(1,z_{L}(1)-z_{L}(0)),\qquad S^{1}(X^{R})=(1,z_{R}(1)-z_{R}(0)).$
+The time-coordinate increments cancel in the adjacent-window distance, so the depth-1 level-aware signature distance is
+$\|S^{1}(X^{L})-S^{1}(X^{R})\|_{2}=|(z_{L}(1)-z_{L}(0))-(z_{R}(1)-z_{R}(0))|.$
+Thus, for affine windows $z_{L}(u)=a_{L}+b_{L}u$ and $z_{R}(u)=a_{R}+b_{R}u$, the depth-1 signature distance reduces to $|b_{L}-b_{R}|$: it compares endpoint slopes over the two windows, not their mean level difference.
+The fatigue-specific rule in Algorithm 1 therefore combines a signature displacement or shape-change statistic with a separate directional mean-drop gate,
+$G_{\tau}=\bar{z}_{L(\tau)}-\bar{z}_{R(\tau)}.$
+This gate is close in spirit to adjacent-window mean-drop or MOSUM-style summaries, while the higher signature levels add terms that can respond to ordering and curvature within the windows.
+If one instead augments the path with a cumulative-level coordinate,
+$Q(u)=\int_{0}^{u}z(v)\,dv,$
+then the first-level displacement in $Q$ equals the window mean over $[0,1]$.
+In that alternative embedding, the depth-1 adjacent-window distance directly contains $|\bar{z}_{L}-\bar{z}_{R}|$.
+This shows that simple mean-drop detectors are recoverable as low-order signature features under an appropriate embedding, and it clarifies the incremental role of the higher-order path terms evaluated in the benchmark.
+C.3 Detection power intuition
+Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Suppose the process has different distributions before and after $\tau$:
 $Y_{t}\sim\begin{cases}\mathcal{F}_{1}&\text{if }t<\tau\\ \mathcal{F}_{2}&\text{if }t\geq\tau\end{cases}$
-For window size $w$ and threshold $\theta=\mu_{D}+k\sigma_{D}$, the detection power (probability of correctly identifying the change) is:
-$\text{Power}=P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{w\cdot\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}{2\sigma^{2}}\right)$
-where $\text{KL}(\cdot||\cdot)$ is the Kullback-Leibler divergence and $\sigma^{2}$ is the variance of the signature distance under the null hypothesis.
-Proof.
-Under $H_{1}$ (change present), windows $W_{1}$ and $W_{2}$ straddling the change point contain samples from different distributions.
-The signature distance $D$ between these windows can be bounded using the data processing inequality:
-$D^{2}=\|S(X_{W_{1}})-S(X_{W_{2}})\|^{2}\geq w\cdot d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})^{2}$
-where $d_{\text{TV}}$ is the total variation distance.
-By Pinsker’s inequality:
-$d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})\geq\sqrt{\frac{1}{2}\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}$
-Under $H_{0}$, the signature distances concentrate around their mean; a complete derivation depends on dependence assumptions in the time series and on the distribution of the signature estimator.
-Using concentration-style reasoning:
-$P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{(\mathbb{E}[D|H_{1}]-\theta)^{2}}{2\sigma^{2}}\right)$
-Substituting lower bounds on $\mathbb{E}[D|H_{1}]$ yields the stated qualitative dependence.
-The intent is to highlight that detection improves with larger window sizes and with greater separation between pre- and post-change regimes.
-∎
-C.3 Convergence Properties as Window Size Increases
-Theorem 3 (Asymptotic consistency (informal)).
-Let $w$ be the window size and $n$ be the total number of observations.
-As $w\to\infty$ with $w=o(n)$, the signature-based change point estimator $\hat{\tau}_{w}$ converges to the true change point $\tau$ in probability:
-$P(|\hat{\tau}_{w}-\tau|>\epsilon)\to 0\quad\text{as }w\to\infty$
-for any $\epsilon>0$.
-Proof.
+Under $H_{1}$ (change present), adjacent windows straddling the change point may have different expected path shapes, and therefore larger expected signature distances, than windows drawn from a stable regime.
+Under $H_{0}$, the distribution of signature distances depends on the sampling process, window overlap, exposure levels, and serial dependence.
+A complete power analysis would therefore require a specified stochastic model for impressions, clicks, and latent CTR dynamics.
+We do not claim a closed-form lower bound from KL divergence or total variation to a realised Euclidean signature distance; data-processing inequalities concern divergences between distributions after measurable transformations, not deterministic lower bounds on one sampled path.
+Standard Pinsker’s inequality is
+$d_{\mathrm{TV}}(P,Q)\leq\sqrt{\frac{1}{2}\mathrm{KL}(P\|Q)},$
+or equivalently $\mathrm{KL}(P\|Q)\geq 2d_{\mathrm{TV}}(P,Q)^{2}$.
+This inequality can be useful in a fully specified distributional analysis, but it is not sufficient by itself to prove a lower bound for the detector used here.
+C.4 Convergence Properties as Window Size Increases
+Let $w$ be the window size and $n$ be the total number of observations. A plausible consistency target for a window-based estimator is localisation at the scale of the window, for example:
+$P(|\hat{\tau}_{w}-\tau|/w>\epsilon)\to 0\quad\text{as }w\to\infty,\quad w=o(n).$
+This scaled statement is weaker and more appropriate than exact recovery of an integer changepoint index.
 Define the localised signature distance process:
 $D_{t}(w)=\|S(X_{[t-w,t]})-S(X_{[t,t+w]})\|$
-Under regularity conditions (e.g., bounded $p$-variation and appropriate mixing), this process satisfies:
-  1. 1.
-    For $t$ far from $\tau$: $\mathbb{E}[D_{t}(w)]=O(w^{-1/2})$ (by CLT for signatures)
-  2. 2.
-    For $|t-\tau|<w$: $\mathbb{E}[D_{t}(w)]\geq c\cdot\min(|t-\tau|/w,1)$ for some $c>0$
-The change point estimator is:
-$\hat{\tau}_{w}=\arg\max_{w<t<n-w}D_{t}(w)$
-By the continuous mapping theorem and the fact that $D_{t}(w)$ has a unique maximum near $\tau$ (with high probability), we have:
-$\hat{\tau}_{w}\to\tau\quad\text{in probability}$
-The rate of convergence depends on additional assumptions and is stated here for intuition only. Larger window sizes provide more stable estimation at the cost of reduced temporal resolution. ∎
-C.4 Computational Complexity Analysis
-Lemma 1 (Signature computation complexity (order of growth)).
+Under regularity conditions (e.g., bounded $p$-variation, appropriate mixing, and a unique separation of pre- and post-change regimes), one would need to show that this process is maximised near $\tau$ with high probability.
+A proof would also need to account for overlapping windows and threshold estimation.
+The present paper does not rely on such a theorem; it uses simulation evidence and reports alert burden explicitly.
+C.5 Noise and exposure considerations
+The signature is a continuous functional of the path under appropriate rough path metrics, so small perturbations of a path produce controlled perturbations of its signature.
+This continuity supports the use of signatures as stable path features.
+However, observed CTR has heteroskedastic binomial noise:
+$\operatorname{Var}(C_{t}/I_{t}\mid I_{t},p_{t})=\frac{p_{t}(1-p_{t})}{I_{t}}.$
+Consequently, min-max scaling of realised CTR paths does not remove the dependence of noise level on impression volume.
+Exposure-aware variants–for example signatures of $(t,\widehat{p}_{t},\log I_{t})$, distance weights based on impressions, or detectors applied to the same Beta–Binomial smoothed series used for the evaluation target–are natural extensions and should be evaluated before making strong robustness claims.
+C.6 Computational Complexity Analysis
+Lemma 1 (Per-window signature computation complexity (order of growth)).
 For a path of length $w$ in $\mathbb{R}^{m}$ and truncation depth $d$, the number of truncated signature terms grows as $\sum_{k=1}^{d}m^{k}$. A naive computation scales approximately as:
-$\mathcal{O}(w\cdot m^{d}\cdot d).$
-In practice, log-signature representations reduce constant factors substantially; for fixed $m=2$ and small $d$ (as used in our experiments), the overall detection pipeline is effectively linear in the number of time points.
+$\mathcal{O}\!\left(w\sum_{k=1}^{d}m^{k}\right),$
+which is $\mathcal{O}(wm^{d})$ for fixed $m>1$ up to constants.
+In practice, log-signature representations can reduce constant factors.
+Across $N$ admissible window positions, the detector has order $\mathcal{O}(Nw\sum_{k=1}^{d}m^{k})$.
+It is effectively linear in $N$ only when the window size $w$, path dimension $m$, and truncation depth $d$ are treated as fixed constants.
 Proof.
-The dominant cost is evaluating iterated integrals (or equivalent algebraic updates) over $w$ increments.
-Since the number of terms grows exponentially in $d$, practical implementations fix $d$ (e.g., $d\in\{3,4\}$) and treat it as a small constant.
+For one window, the dominant cost is evaluating iterated integrals (or equivalent algebraic updates) over $w$ increments for the retained signature coordinates.
+The stated order assumes a naive per-increment update in which each retained coordinate is updated once with constant-time algebraic work, so the cost per increment is proportional to $\sum_{k=1}^{d}m^{k}$.
+Since the number of terms grows exponentially in $d$ and polynomially in $m$ through $\sum_{k=1}^{d}m^{k}$, practical implementations fix both $m$ and $d$ (e.g., $m\in\{2,3\}$ and $d\in\{3,4\}$) and treat them as small constants.
+Computing the Euclidean distance between two signatures is also linear in the retained feature dimension and becomes constant per window only after $m$ and $d$ are fixed.
+The full detector computes such quantities across $\mathcal{O}(T)$ window positions when $w$, $m$, and $d$ are fixed, followed by linear-time thresholding and post-processing steps.
 ∎
-C.5 Robustness to Noise and Outliers
-Theorem 4 (Noise robustness (informal)).
-Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$.
-Then the expected signature distance satisfies:
-$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right),$
-where $C$ depends on the truncation depth $d$ and regularity of $X$.
-Proof.
-The signature is a continuous functional of the path. By the rough path stability theorem, small perturbations in the path lead to controlled changes in the signature. For additive noise:
-$S(\tilde{X})-S(X)=S(X+\epsilon)-S(X)=\sum_{k=1}^{d}\int\cdots\int d\epsilon_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}+O(\|\epsilon\|^{2})$
-Taking expectations and using the independence of noise:
-$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]=\sum_{k=1}^{d}\mathbb{E}\left[\left\|\int\cdots\int d\epsilon_{t_{1}}\otimes dX_{t_{2}}\otimes\cdots\right\|^{2}\right]$
-By Itô isometry and the bounded variation of $X$:
-$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq\sigma^{2}_{\epsilon}\cdot w\cdot\sum_{k=1}^{d}\frac{\text{Var}(X)^{k-1}}{k!}$
-The sum converges to a finite constant under bounded-variation assumptions, giving the stated qualitative dependence.
-This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness.
-∎